### v1

# A Portable and Stealthy Inaudible Voice Attack Based on Acoustic Metamaterials

Abstract
We present MetaAttack, the first approach to leverage acoustic metamaterials for inaudible attacks for voice control systems. Compared to the state-of-the-art inaudible attacks requiring complex and large speaker setups, MetaAttack achieves a longer attacking range and higher accuracy using a compact, portable device small enough to be put into a carry bag. These improvements in portability and stealth have led to the practical applicability of inaudible attacks and their adaptation to a wider range of scenarios. We demonstrate how the recent advancement in metamaterials can be utilized to design a voice attack system with carefully selected implementation parameters and commercial off-the-shelf components. We showcase that MetaAttack can be used to launch inaudible attacks for representative voice-controlled personal assistants, including Siri, Alexa, Google Assistant, XiaoAI, and Xiaoyi. The average word accuracy of all assistants is 76%, with a range of 8.85 m.
Index Terms:
Inaudible Attack, Acoustic Metamaterials, Voice Control Systems, Security and Privacy.
publicationid: pubid: $0$–$00$ 2021 IEEE
I INTRODUCTION
Voice control is a natural way for human-computer interactions, commonly found in mobile and smart home applications [1, 2, 3, 4, 5, 6]. Given its capacity to control a broad spectrum of tasks – from making phone calls to managing devices like surveillance cameras and even initializing bank transfers – voice control systems have increasingly become the target of attacks [7, 8, 9, 10, 11, 12, 13, 6].
Recent research has demonstrated that the voice control mechanisms in mainstream personal assistant systems, like Amazon’s Alexa and Apple’s Siri, are susceptible to inaudible attacks [7, 8, 9]. Such attacks modulate voice commands onto ultrasonic carriers, generating commands imperceptible to human ears but can be accepted by voice control systems. Despite some levels of success in specific scenarios, existing attacking methods have a short attack range of up to 3 m. However, a close-in attack within 3 m can be easily detected by many vigilant individuals [14]. Extending this range by simply increasing transmission power would lead to audible noise, thereby compromising the stealthiness of the attack. More recent work, as represented by LipRead [14], divides the spectrum of the inaudible commands to reduce audible leakage and utilizes loudspeaker arrays to expand the attack range. LipRead is shown to be able to launch inaudible attacks from distances up to 7.62 m using 61 loudspeakers. Another representative system, DolphinAttack [15], can reach 19.8 m for one model of the device (iPhone X), requires 40 speakers, and uses an external power supply and amplification equipment. However, the division of the spectrum of inaudible commands reduces the accuracy of inaudible commands, and the size of the amplification equipment makes it difficult to launch the attack in many practical public scenarios.
This paper demonstrates the feasibility of executing long-range, inaudible attacks through a compact audio modulation system that can easily fit into a modest-sized travel bag. We introduce MetaAttack, the first inaudible attack that leverages the emergent acoustic metamaterials technology [16, 17, 18, 19, 20, 21]. Acoustic metamaterials are engineered to enable the manipulation, control, and transformation of acoustic waves in ways impossible to conventional materials. These metamaterials’ dimensions are smaller than the wavelengths of the sound waves they interact with. By carefully designing the elements’ geometry, size, shape, and spatial arrangement, MetaAttack can precisely control the acoustic wave propagation via compact devices [22, 23, 24]. For example, Fig. 1 depicts a possible attack scenario where MetaAttack is covertly placed in a public place next to a window to target a device located in another building within 8.5 m.
Figure 1: A possible attacking scenario of MetaAttack where an attacker can launch an inaudible attack in a public place to remotely control the victim device in another building. In this scenario, the attack device must be stealthy and does not require the use of other continuously powered devices such as DC power supply.
While acoustic metamaterials are shown to be useful in other domains [25, 26, 27, 28], there is no prior work for using them to build a practical inaudible attack system. In our work, we introduced two changes to improve the metamaterial’s functionality. First, we refined the internal structure to expand its filtering capabilities, thus minimizing audible leakage during inaudible attacks. Second, we converted the metamaterial into a piston sound source  [29, 30] by carefully designing the incident sound wave’s shape, increasing the energy intensity of inaudible commands, and extending the attack range. We call our enhanced design a multi-functional acoustic metamaterial (MAM), as it not only filters out audible leakage but also amplifies the energy of inaudible commands. With these advancements and a compact design, MetaAttack can execute long-range inaudible attacks while remaining highly concealed.
We showcased that MetaAttack can be implemented with low-cost, commercial, off-the-shelf components like Raspberry Pi, amplifiers, loudspeakers, servo, and a 3D-printed MAM. We thoroughly evaluate MetaAttack under different practical scenarios under different attack ranges, tasks, noise levels, obstacles, and targeting different victim devices. We compare MetaAttack against three prior inaudible attacks [8, 9, 14]. Experimental results show that MetaAttack is robust under different evaluation scenarios and tasks and can successfully launch an audible attack within a 8.85 m range with a noise level of up to 65 dB. Compared with prior attacking methods, MetaAttack achieves the good accuracy and attack range (76% average accuracy under a 8.85 m attack range) with the smallest sized device.
This paper makes the following contributions:
  • We present the first acoustic metamaterial-based inaudible attacks and released a complete process demonstration of MetaAttack [31, 32] and open-sourced the code [33];
  • We demonstrate how acoustic metamaterial can be repurposed to improve the capability of inaudible attacks;
  • We showcase how a practical attacking device can be built from commercial off-the-shelf components and discuss the potential countermeasures.
II BACKGROUND
II-A Inaudible Attacks
Inaudible attacks leverage ultrasonic frequencies (sounds above 20 kHz), which the microphones in smart devices can often detect even though humans cannot hear them. The concept behind these attacks is to issue commands to voice-activated devices without the device owner’s knowledge.
Specifically, when microphones used in voice control systems receive input signals exceeding 20 kHz, due to the non-linearity effect, the microphones output additional non-linear signals ${S_{in}}^{2}(t)$ [15]. ${S_{in}}(t)$ as shown in Eq.1:
${S_{in}}(t)={V_{attack}}\left(t\right)\cos\left({2\pi{f_{carrier}}t}\right)+\cos\left({2\pi{f_{carrier}}t}\right)$ (1)
where ${V_{attack}}\left(t\right)$ is the voice control signal, and $\cos\left({2\pi{f_{carrier}}t}\right)$ is the high frequency carrier signal. When the attack signal is input to the microphones, the microphones output the signal $S_{attack}(t)$ as shown in Eq.2:
$S_{attack}(t)$ $=S_{h}+\frac{A_{2}}{2}\left(1+V_{attack}^{2}(t)+2V_{attack}(t)\right)$ (2)
where ${V_{attack}}\left(t\right)$ can be recognized by the microphones, allowing the inaudible attacks to be successfully realized.
By utilizing the modulation method described above, an adversary encodes commands in ultrasonic frequencies that the microphones of nearby voice-activated devices can pick up. For example, these commands could be “turn off the surveillance camera” if the system is connected to smart home security, “make a purchase” and “transfer money” if linked to financial services, “sending counterfeit messages”, or “calling number X” for eavesdropping during a meeting. The ultrasonic commands are transmitted over the air, potentially from a loudspeaker system or a specialized device, and received by the target device. Because these sounds are inaudible to humans, the device owner remains unaware that their device is being controlled remotely. However, the voice recognition of the targeted device can process these sounds as normal voice commands and execute the action without the user’s knowledge or consent, posing a significant vulnerability for voice-controlled smart devices.
The effectiveness of inaudible attacks is compromised by the inherent limitation of a limited attack range compared to laser [34] and wireless signal attacks [35], typically ten feet [8, 9]. This limitation results from the audible leakage at the point of transmission during inaudible attacks. Concealment requires a reduction in output power [14], thereby limiting the range of inaudible attacks. As a result, the impact of an inaudible attack decreases significantly as the distance between the sonic emitter and the target increases, making it less suitable for long-range targeting than other forms of attack.
II-B Threat Model
We focus on implementing an inaudible acoustic attack that is portable, miniaturized, and out of detection range. These characteristics address the need for attacks in portable and stealthy environments. Our attack does not require the attacker to access the victim’s device physically. However, we assume the attacker can synthesize the target’s voice using, e.g., deep-learning-based voice generation from a few training samples [8]. We believe this assumption is reasonable in many day-to-day scenarios where the victim’s voice samples can be collected from public places, a multi-user meeting, or audio and video clips available to the attacker. We also assume that the victim was not intact with the device during the attack and surroundings are not completely enclosed so that it is possible to transmit ultrasounds.
II-C Practical Limitations of Prior Methods
Launching inaudible voice commands involves modulating the control voice signal. However, nonlinear effects from devices like loudspeakers can cause audible leakage, particularly when operating at high power levels, to extend the attack range. Previous works [8, 9] limited device power to prevent sound leakage, but doing so limits the attacking range to around 3 m. To enable longer-range inaudible attacks, LipRead [14] segments the inaudible command spectrum into low-frequency parts that are less detectable by humans. While representing an improvement, LipRead also has drawbacks. It requires an array of 61 loudspeakers to reach up to an attacking distance of 7.62 m, increasing the risk of detection, and the method of segmenting the spectrum may lower command accuracy due to potential recombination or interference during transmission. Similarly, although Yan et al. [15] were able to extend the ultrasonic attack range to 19.8 m, the attack needed a signal generator and a constant DC power supply, which makes it less practical for portable mobile attacks. Additionally, the use of 40 loudspeakers was necessary for the attack device. To more clearly illustrate the limitations of existing studies regarding the distance of the inaudible sound attack, we present a visual representation in Table I. Most of the previous research has been on defeating voiceprint recognition [14, 7, 9, 8], but these studies are difficult to conduct in real-world scenarios. Some of the work requires close proximity to the attacking device, and although long-range attacks can be achieved, they require the introduction of more auxiliary devices [15, 14, 7] that are not stealthy. A viable attack therefore requires a balance between lightness and range.
TABLE I: Comparison of prior inaudible attack research in terms of attack range, portability and stealthy
System name System setup DC power supply Signal generator Speaker number Attack range Portability Stealthy
DolphinAttack (Long-range System) [15] [Uncaptioned image] Yes (Approximately 11 MacBook Pro) Yes (Approximately 12 MacBook Pro) 40 19.8m (iPhoneX) 10m (iPhoneSE) No No
LipRead [14] [Uncaptioned image] Yes Yes (Approximately 7 MacBook Pro) 61 7.62m No No
BackDoor [7] [Uncaptioned image] Yes Yes (Approximately 7 MacBook Pro) 18 3.5m No No
EchoAttack [9] [Uncaptioned image] No Yes (Approximately 7 MacBook Pro) 40 2.23m Yes Yes
VRIFLE [36] [Uncaptioned image] No No 8 1.8m Yes Yes
DolphinAttack (Portable System) [8] [Uncaptioned image] No No 1 0.27m Yes Yes
II-D Motivations
Most of the attack scenarios presented in previous studies are based on the assumption that the device to be attacked is stationary in a specific location [14, 7, 15, 8], or that a large speaker array is used if the distance is too great [14, 15] - assumptions that are not always true. This is because in many public environments, using external power through amplification devices could easily expose the malicious intent behind an attack, often leading to early detection before it is initiated. As shown in Table I, if the attacking system lacks DC power support, its attack range is limited[14, 15]. In this paper, for parts A and B of Fig 2, we aim to design an attack device that is portable and concealable, without the need for an external DC auxiliary power supply, and that can be implemented over long distances.
Figure 2: Attack Range/Device Size quadrant chart. (Large: with DC power supply, signal generator and large speaker array. Medium: with a signal generator and large speaker array. Small: without above devices.)
II-E Metamaterials
We present the first inaudible attack using acoustic metamaterials. Our key insight is that a carefully engineered metamaterial can be designed with structures that interact with sound waves in specific ways, allowing them to block or absorb sound at certain frequencies while letting other frequencies pass through [37, 38]. This means we can employ metamaterials to create filters that target specific sound frequencies while allowing airflow, enhancing the power of inaudible commands and minimizing audible leakage. In this work, we leverage metamaterials to design a compact, passive device, offering two main benefits over existing inaudible attacking methods. First, the MAM preserves the accuracy of inaudible commands by filtering audible sounds without splitting their spectrum. Secondly, it enables using a few loudspeakers to replicate the effect of a larger array, facilitating covert and long-range attacks.
III OVERVIEW of OUR APPROACH
As shown in Fig. 3, MetaAttack is a 3-step inaudible attack, described in detail below.
Figure 3: Overview of MetaAttack with three steps: Processing inaudible commands, Attacking victim’s device, and Activating feedback mechanism.
Processing inaudible commands. The initial phase of our attack is to process and modulate voice commands. In this work, we first captured the target’s voice and synthesize the target’s voiceprint (see Section II-B). Then we modulated the voice command groups generated by Google Text-to-Speech AI [39] and voiceprints onto ultrasound carrier waves using the modulation method discussed in Section II to form inaudible commands [14]. Each voice command group used for the attack contains three types of inaudible instructions: commands for muting the victim devices, commands for executing attacks, and commands for feedback. For instance, mute commands might include “speak 6% ($MuteCmd$)” for Siri or “turn the volume to 1 ($MuteCmd$)” for other devices [35, 40], attack commands might include “send message to person X ($AttackCmd$),” while feedback commands could be “turn on hotspot ($FeedbackCmd_{1}$)” and “turn off hotspot ($FeedbackCmd_{2}$)”. Subsequently, a group of inaudible commands is transmitted through a loudspeaker array. During this process, any audible leakage is filtered out, and our MAM amplifies the inaudible commands.
Attacking victim’s device. MetaAttack uses beamforming to send the inaudible commands, but this requires the signal to be transmitted within a specific angle of approximately 12°. To position the signal transmission angle, we use a servo to drive MetaAttack to point to the victim device within a 180° range in steps of 12°. We first send $MuteCmd$ to lower the victim device’s loudspeakers to an inaudible level[35, 40]. Then, we send $AttackCmd$ to achieve the purpose of the attack. Finally, we send $FeedbackCmd_{1}$ and $FeedbackCmd_{2}$ to activate the feedback mechanism.
Activating feedback mechanism. The feedback mechanism uses the Raspberry Pi to detect the state changes caused by the wireless signal switching command received by the target device, thereby evaluating whether the attack is successful. If the hotspot information of the device completes state transitions between appearance and disappearance in the detection list of MetaAttack after the attacker sends a command group, we consider the execution of the command group successful. Considering that the SSID list within the attack range is not static, the SSID information from weak wireless signals of non-target devices may randomly appear or disappear, we use iwlist [41] to monitor the SSID and signal strength of wireless signals in real-time. Specifically, when the SSID information from weak wireless signals appears or disappears, its signal strength typically shows a gradual increase or decrease. In contrast, the wireless signal from the victim device triggered by $FeedbackCmd$ does not undergo such gradual changes when it appears or disappears. Therefore, after sending $FeedbackCmd$, we use iwlist to identify signals that appear or disappear directly, thereby eliminating the interference from weak wireless signals in our assessment.
As shown in Algorithm 1 and Fig. 4, we determine the attack’s success by identifying the hotspots of smartphones or the Bluetooth connectivity of smart home devices. To illustrate the feedback mechanism for smartphones, we break it down into four sequential steps:
Step 1. After completing the transmission of the $MuteCmd$ and $AttackCmd$ (line 1), the Raspberry Pi uses iwlist to record the identified hotspot devices and stores the information in the L1 list (line 1). The system then proceeds to send the $FeedbackCmd_{1}$ to store the records in the L2 list (line  1).
Step 2. The system continues to send $FeedbackCmd_{2}$ and stores the received hotspot information in list L3 (line  1). If there are discrepancies in the hotspot signals ($dif_{1}$) between the L2 and L3 lists, the process moves to Step 3 (lines  11). Otherwise, the MetaAttack is rotated to the next angle (line  1).
Step 3. MetaAttack will sequentially send $FeedbackCmd_{1}$ and then instruct the Raspberry Pi to retrieve the L4 hotspot list (line  1). If there are differences in the hotspot signals ($dif_{2}$) between the L3 and L4 lists, and the intersection of $dif_{1}$ and $dif_{2}$ contains only one device, a successful stealth attack is detected (lines  11). If not, the MetaAttack will be adjusted to the next angle (line  1).
Step 4. To maintain the stealth of the attack, this step focuses on restoring the victim to its original state before the attack (lines  11).
After these four steps, the system’s attack feedback mechanism ensures the success rate of the inaudible attack. To confirm the feasibility of our attack process, we have released a complete process demonstration [31] and open-sourced the feedback code [33].
Figure 4: Feedback overview of MetaAttack (After using $MuteCmd$ and $AttackCmd$ to achieve the purpose of inaudible attacks, using $FeedbackCmd_{1}$ and $FeedbackCmd_{2}$ to provide feedback)
Input: $AttackCmd,\ FeedbackCmd_{1},\ FeedbackCmd_{2}$
[…]
2       $Send\ MuteCmd\ and\ AttackCmd$
3       $Activate\ iwlist$
4       feedbackMechanism()
5       if $\textnormal{{feedbackMechanism(}}\textnormal{\emph{}}\textnormal{{)}}\ \textbf{==}\ true$  then
6             break
7       end if
8
9 end foreach
10
11Function feedbackMechanism():
       // $Achieve\ feedback\ mechanism$
12       Step1()
13       Step2($L_{1},L_{2}$)
14       return  Step2($L_{1},L_{2}$)
15
16
17
18Function Step1():
       // $Attempt\ to\ turn\ on\ victim^{\prime}s\ hotspot$
19       $Receive\ hotspot\ list\ L_{1}$
20       $Send\ FeedbackCmd_{1}\ and\ receive\ hotspot\ list\ L_{2}$
21       $\textbf{return}\ L_{1},L_{2}$
22
23Function Step2($L_{1},L_{2}$):
       // $If\ any\ hotspot\ signal\ disappears,\ moves\ to\ Step3$
24       $Send\ FeedbackCmd_{2}\ and\ receive\ hotspot\ list\ L_{3}$
25       $dif_{1}=L_{2}-L_{3}$
26       if $dif_{1}.length\neq 0$ then
27             Step3($L_{1},L_{3},dif_{1}$)
28             return  Step3($L_{1},L_{3},dif_{1}$)
29      else
30             $\textbf{return}\ false$
31       end if
32
33
34
35Function Step3($L_{1},L_{3},dif_{1}$):
       // $If\ hotspot\ signal\ reappears,\ attack\ is\ successful$
36       $Send\ FeedbackCmd_{1}\ and\ receive\ hotspot\ list\ L_{4}$
37       $dif_{2}=L_{4}-L_{3}$
38       $Target\ list=dif_{1}\cap dif_{2}$
39       if $Target\ list.length==1$ then
40             Step4($Target\ list,L_{4},L_{1}$)
41             $\textbf{return}\ true$
42      else
43             $\textbf{return}\ false$
44
45       end if
46
47
48Function Step4($Target\ list,L_{4},L_{1}$):
       // $Restore\ the\ hotspot\ status\ of\ the\ victim\ device$
49       if $L_{4}-L_{1}==Target\ list$ then
50             $Send\ FeedbackCmd_{2}$
51
52       end if
53
54
Algorithm 1 The framework of attack overview
IV Metamaterial Design for Covert Long-Range Attack
IV-A Design Goals and Challenges
To implement a covert long-range inaudible attack system, we need to achieve two goals:
Design Goal 1. Filter out audible leakage. Maintaining the stealth nature of the attack requires filtering out audible leakage in the 100 Hz - 4000 Hz frequency range since the energy of audible leakage is concentrated in this spectrum. This filtering process is critical to maintaining the stealth nature of inaudible attacks and preventing their detection by acoustic leakage.
Design Goal 2. Increase the range of inaudible commands. To further achieve long-range attacks, we need to increase the beamforming effect of the inaudible acoustic attack command. This is because when beamforming is improved, the sound waves can be more precisely focused in the direction of a specific victim device, reducing energy dispersion and loss, which allows the sound waves to travel farther.
To achieve these two design goals, the system design must address two challenging issues.
Design Challenge 1. How to design an acoustic metamaterial with a fixed geometrical structure, which ensures that the sound signal is not leaked (for Design Goal 1) but also allows the inaudible sound signal to propagate over a longer distance (for Design Goal 2).
Design Challenge 2. How to achieve a lightweight and portable design that satisfies the stealthy nature of the attack implementation process(for Design Goal 1). This design requires only a minimum number of speakers and an optimal arrangement.
IV-B Filter Out Audible Leakage
To address Challenge 1, we introduce a novel Leak-shielding Acoustic Metamaterial (LAM). The basic filtering mechanism of the LAM is shown in Fig. 5a. Its internal structure features a circular hole-like design (with a diameter d of 22.5 mm) surrounded by a spiral path (with a width D of 100 mm and a pitch P of 34.542 mm) [25]. When sound waves of certain frequencies pass through the metamaterial, the path variances induced by its internal structure (e.g., path 1 and path 2) result in phase opposition, leading to filtration by reflection [42, 43, 44, 45]. However, since the current structure generates a limited number of paths, it can only provide narrowband filtering (600 Hz - 1900 Hz). This limits its effectiveness in eliminating audible leakage to 100 Hz - 4000 Hz. To effectively counteract audible leakage, our LAM has undergone significant development:
Expanding the filtering frequency range to 100 Hz - 4000 Hz. To further expand the LAM’s filtering range, we leverage a key observation in existing works: an opening at the connection between the circular hole-like structure and the surrounding spiral path can expand the LAM’s filtering range [25]. The reason for this observation is that an opening creates more paths, allowing more frequencies to be filtered out through phase opposition. One question that naturally arises is how to determine the opening geometric size. Through our simulation in COMSOL, we find that the height (h) and the angle ($\gamma$) of the opening both affect the filtering range of LAM. Therefore, we have tried various combinations of h and $\gamma$, then as shown in Fig. 5b, we find the opening height $h=3~{}mm$ and angle $\gamma={68.7^{\circ}}$ can expand the filtering range to 100 Hz - 4000 Hz.
Figure 5: (a) The filtering principle of LAM. (b) LAM- Leak-shielding acoustic metamaterial
Even though our LAM can utilize its unique structure to filter out audible leakage in front of the loudspeaker array through phase opposite (as shown in Fig. 5a), it cannot prevent audible leakage from leaking from the sides or back of the loudspeaker array. To address this issue, our LAM has undergone another development:
Preventing the audible leakage from leaking to other directions. As shown in Fig. 5b. Firstly, to prevent audible leakage from leaking from the sides of the loudspeaker array, we increase the thickness (${L_{1}}$) and the length (${L_{3}}$) of the wall on both sides of the LAM ($H=50mm$) to wrap around the loudspeaker array. This causes audible leakage on both sides to be reflected by the wall. Secondly, to eliminate the audible leakage on the back, we have added a cover plate at the rear of the LAM and filled the front of the cover plate with melamine sponge (${L_{2}}$) so that the audible leakage is eliminated by the thermal viscosity loss caused by the melamine sponge. Where length ${L_{3}}=17.5~{}mm$ and its value equals the thickness of the loudspeaker array.
To investigate the effects of different wall thicknesses $L_{1}$ and melamine sponge thickness $L_{2}$ on eliminating sides and back audible leakage, we simulated them in COMSOL. Since the vast majority of voice commands have the energy to 100 Hz - 4000 Hz, we set the frequency in the range of these commands. Then, we let the sound waves emitted by the sound source pass through the LAM and calculated the energy of the sound waves on the sides and back of the LAM, measured in sound pressure level (SPL). The simulation validation results of some representative $L_{1}$ and $L_{2}$ are shown in Figure 6a and Figure 6b. To ensure that the audible leakage remains inaudible to people, prior works, such as LipRead [14], have demonstrated that the “Leakage Threshold”, should be 5 dB below the “Threshold of Hearing Curve”. As a result, we derive the “Leakage Threshold” by subtracting 5 dB from the “Threshold of Hearing Curve”, as depicted by the red line in Figure 6. The results demonstrate that when the wall thickness $L_{1}$ is greater than 2 mm and the melamine sponge thickness $L_{2}$ is greater than 40 mm, audible leakage is below the “Leakage Threshold”. Therefore, on the premise of considering the filtration efficiency and compactness of LAM, we have determined the wall thickness ${L_{1}}=2~{}mm$ and the melamine sponge thickness ${L_{2}}=40~{}mm$. The total length of the LAM is ${L_{4}=H+L_{1}+L_{2}+L_{3}}=109.5~{}mm$.
Figure 6: (a) Filtering effect on sides of LAM. (b) Filtering effect back of LAM. (c) Filtering effect in front of LAM.
To verify the filtering effect of the LAM, we tested its filtering performance in COMSOL. The simulation validation results are shown in Figure 6c. The result demonstrates that the audible leakage after passing through the LAM filter is below the “Leakage Threshold”, indicating the effectiveness of our LAM in mitigating audible leakage.
Filtering audible leakage against inaudible attacks only addresses the stealth of the attack. Therefore, additional design improvements are required to further enable longer-range attacks with a reduced number of loudspeakers. The key to this advancement is to keep the sound from leaking while ensuring sufficient range.
IV-C Increase the Range of Inaudible Commands
In the above improvements, the LAM can only play a role in filtering audible leakage in inaudible attacks. However, by carefully designing the shape of incident sound waves, the LAM can be transformed into a piston sound source. The piston sound source can enhance the high-frequency sound waves through acoustic resonance [29], allowing a small loudspeaker array to perform the same beamforming effect as a large loudspeaker array to enhance inaudible commands.
Figure 7: Enhancement effect of inaudible commands in LAM
Figure 8: MAM (Multi-functional Acoustic Metamaterial ($H_{2}=127.5~{}mm$).
As shown in Figure 8, LAM consists mainly of a barrier and a circular piston in the center, and its structure is similar to the piston sound source [29, 30]. When the inaudible commands pass through the LAM, the circular barrier causes them to converge into the circular hole-like structure, thus enhancing them.
However, from the simulation results, we found that the surrounding spiral path inside LAM would disrupt the incident plane wave of inaudible commands, failing to form a piston sound source (the prerequisite for transforming the LAM to a piston sound source is that the incident waveform of inaudible commands is a plane wave [46, 47]).
TABLE II: The impact of various factors on the enhancement effect of inaudible commands
Number of Loudspeakers Range (R) Carrier Frequency Maximum SPL
2 22 mm 40500 Hz 128 dB
4 14 mm 40000 Hz 130 dB
6 14 mm 40000 Hz 134 dB
8 11 mm 39500 Hz 135 dB
10 15 mm 40500 Hz 138 dB
12 18 mm 40200 Hz 142 dB
14 18 mm 40200 Hz 142 dB
16 18 mm 40200 Hz 142 dB
Figure 9: The sound field of loudspeaker array
Figure 10: The results of enhancement effect.
Figure 11: The prototype of MetaAttack
Figure 12: The covert MetaAttack
To solve this issue, we let the inaudible commands pass through a LAM under different conditions and observed its waveform. Then, we find that whether the inaudible commands can form plane waves depends on three factors: (1) the number of loudspeakers; (2) the range between the loudspeaker array and circular hole-like structure; (3) carrier frequency of inaudible commands. To investigate the impact of these factors, we simulated them in COMSOL. After simulating these factors, we obtained the max SPL under different conditions in Table II.
We traverse all the possible combinations of the number of speakers, range (R), and carrier frequency to investigate their impact. Table II provides the optimal combinations of range (R) and carrier frequency under different numbers of loudspeakers. The highest gain is achieved when 12 loudspeakers are incident at 40200 Hz, positioned at a range of 18 mm (i.e., $R=18~{}mm$) from the circular hole-like structure. We also found that when using 14 loudspeakers or more, the SPL no longer increases. This is because the 12 loudspeakers perfectly cover the circular hole-like structure in the LAM, and more loudspeakers will be disturbed by other structures. Furthermore, we conducted a comparative analysis of the various layouts of the 12 speakers in the experiment referenced as Section VI-B2. This analysis led to the identification of the optimal configuration, namely, two rows and six columns, which was found to result in the greatest possible attack range.
The final structure is depicted in Fig.  8. This final structure is referred to as the Multi-functional Acoustic Metamaterial (MAM). Its length is ${H_{2}}={L_{4}}+R=127.5~{}mm$. To verify its performance, we first simulate the sound field of the loudspeaker array. The result is shown in Fig. 12. We have proven that under this structure, the waveform of the loudspeakers is a plane wave when it reaches the circular hole-like structure. We then investigate its enhancement effect in COMSOL. As shown in Fig. 12, we find that when inaudible commands meet the above conditions and pass through the MAM, the SPL increases by 11 dB, indicating that the energy of inaudible commands has been enhanced.
V IMPLEMENTATION
V-A MetaAttack Components
The prototype of MetaAttack uses commercial devices and a 3D printed MAM. Fig. 12 shows our prototype system. This metamaterial serves as the basis for MetaAttack, which filters out low-frequency audible sounds and amplifies inaudible commands. The Raspberry Pi acts as a terminal for MetaAttack, facilitating the transmission of commands to attack victims and control the servo. It also detects the hotspot or Bluetooth signal from the victim device to determine the attack’s success. The amplifier amplifies the ultrasonic inaudible commands transmitted by Raspberry Pi. The servo helps MetaAttack to attack over a long distance. A loudspeaker array (12 cm $\times$ 1.4 cm) in MAM transmits inaudible commands with a center frequency of 40 kHz.
V-B Stealthiness and Portability
The stealthiness and portability of inaudible attacks play a significant role in facilitating covert operations, evading detection, targeting diverse devices, and expanding the attack surface to maximize the effectiveness and reach of these malicious activities. As shown in Fig. 12, the length, width, and height of the assembled MetaAttack are 12.75 cm, 11 cm, and 18.8 cm, respectively. Its compact size allows us to carry it conveniently. Therefore, to enhance the stealth capabilities of MetaAttack, we have consolidated all attack devices into a compact travel bag. The bag can be held in hand or slung over the shoulder when performing a silent attack. With its compact size and enhanced stealth features, the MetaAttack in its travel bag configuration is a powerful tool for silent and covert operations.
VI EVALUATION
VI-A Experimental Setup
Target devices. We evaluated four mainstreamed voice-controlled personal assistants on smartphones and smart devices, including Apple Siri [48], Amazon Alexa [49], Google Assistant [50], Xiaomi Xiaoai [51] and Huawei Xiaoyi [52]. In our evaluation, we use twelve devices and the specifications of the devices are given in the Table III.
TABLE III: Victim devices
Manuf. Model OS Voice system
iPhone 14 Pro IOS 17.4.1 Siri
iPhone 13 Pro Max IOS 17.2.1 Siri
iPhone 13 IOS 17.2.1 Siri
Apple iPhone 12 IOS 16.5.1 Siri
Amazon Echo Dot 5th 9698496900h Alexa
Samsung Galaxy S7 Android 8.0.0 Google Assistant
 Google Pixel 3 Android 12.0.0 Google Assistant
Pixel 2 Android 10.0.0 Google Assistant
Redmi K50 Ultra Miui 13.0.10 XiaoAI
 Xiaomi Xiaoai Play 2 1.62.26 XiaoAI
Mi10 Lite Zoom Miui 12.0.6 XiaoAI
 Huawei nova 7 SE 5G HarmonyOS 3.0.0 Xiaoyi
(a) S1
(b) S2
(c) S3
(d) S4
Figure 13: Real attack scenarios: (a) S1: The attacker attacks the victim from another building at a range of 8.5 m, (b) S2: The attacker attacks the walking victim (the victim’s speed is 1 m/s and 1.5 m/s), (c) S3: The attacker sends the inaudible command: “Call 1234567890” to achieve eavesdropping purpose. (d) S4: The attacker attacks the victim in an obstacle environment.
Metrics. Table IV lists the quantified metrics used in our evaluation. The success rate is a key metric for evaluating inaudible attacks. We measure the attack range by checking whether the attack system can achieve a success rate of 50% or more within the range (RSA). Word accuracy measures the ratio at which inaudible commands are correctly recognized. Precision and Recall measure the accuracy of the MetaAttack feedback system.
Experiment design. We conducted a range of experiments, all of which, except experiment S3, used “Send a message to Jack” as the inaudible command to achieve the purpose of sending false information. First, experiments A1 to A5 were designed to evaluate the impact of different factors on MetaAttack. Second, experiments S1 to S4 were conducted to evaluate the attack performance and the accuracy of the feedback system in real-world scenarios. As can be seen from Fig. 13, the suitable attack distances for the S1 - S4 experimental scenarios are all over 4 m, and all of them are situated in outdoor public areas or small private spaces, making them unsuitable for using large external devices. We believe that these are common scenarios, but it is difficult to fully adapt the existing methods, either the stealth requirement is met but the attack distance is not sufficient [7] [9] [36] [8], or the attacking device has a certain range but is not stealthy [14] [15]. Finally, we compared MetaAttack to prior research on inaudible attacks to highlight its superiority. Detailed descriptions of the experiments can be found in Table V.
Our experiments were approved and overseen by the University’s IRB(Institutional Review Board) to ensure compliance with data security measures and ethical guidelines for data processing.
TABLE IV: Evaluating metrics
Metrics Description
Success rate The ratio between the number of successful attacks and the total number of attacks (30 times per experiment) [10, 9].
RSA The Range at which MetaAttack can Successfully Attack (with a success rate of over 50% ) [14].
Word accuracy The ratio of the number of correctly recognized words in inaudible commands to all words [8, 14, 15].
Precision The proportion of successful attacks correctly identified by the feedback mechanism to all successful attacks identified by the feedback mechanism [53, 54, 14].
Recall The proportion of successful attacks correctly identified by the feedback mechanism to all truly successful attacks [53, 54, 14].
TABLE V: List of experimental setting
Test objectives Label Test focus Description
A1 (Sec.VI-B1) Range of attack on different victim devices between MetaAttack and loudspeaker array To test the impact of range and victim devices to MetaAttack, we let MetaAttack attack the various devices (mentioned in Table III) at different ranges.
A2 (Sec.VI-B2) Attack range by different arrangements of loudspeaker arrays To find the optimal arrangement of the 12 speakers (see Sec.IV-C) for MetaAttack, we used six common settings of the speaker array to attack the victim devices.
A3 (Sec.VI-B3) Filtering of audible leakage To test the practical filtering performance of MetaAttack in inaudible attacks, we tested the audible sound leakage in front of and back of MetaAttack at frequencies between 100 Hz - 4000 Hz.
A4 (Sec.VI-B4) MetaAttack’s enhancement functionality To test the practical performance of MetaAttack’s enhancement functionality, we compared the performance of MAM, LAM and loudspeaker array.
Performance under va- rious conditions A5 (Sec.VI-B5) Noise impact To test the impact of background noise on MetaAttack, we continuously increased the noise from 55 dB to 75 dB and tested MetaAttack’s performance in this environment.
Performance in real- world scenarios S1 (Sec.VI-C1) Long-range attacks between high-rise buildings To test MetaAttack’s long-range attack performance in real-world attack scenarios, we attacked the victim devices inside another building, with the two buildings located 8.5 m apart (as shown in Fig. 13a).
S2 (Sec.VI-C2) Attacks against walking victims To test MetaAttack’s performance in attacking against walking victims, we attacked a victim who is moving on the street, and the victim’s speed is 1 m/s and 1.5 m/s (as shown in Fig. 13b).
S3 (Sec.VI-C3) Eavesdropping attacks in indoor environments To test MetaAttack’s eavesdropping performance, we let MetaAttack send “Call 123456780” to attack the victim devices to call the attacker’s number for eavesdropping (as shown in Fig. 13c).
S4 (Sec.VI-C4) Effect of passing individuals on attack range To test MetaAttack’s performance in public environments, we attacked the victim devices in a public room, and there are people along the attack path (as shown in Fig. 13d).
S5 (Sec.VI-C5) Feedback on the results of an inaudible attack To verify whether the feedback mechanism of MetaAttack is accurate, we ran it in S1 - S4 and calculated its Precision and Recall.
Performance compared to prior research Sec. VI-D Advantages of the MetaAttack in terms of concealment To verify that MetaAttack can achieve more covert attacks, we compared the size of its devices with previous attack system.
VI-B Performance under Various Conditions
VI-B1 A1 - Range of attack on different victim devices between MetaAttack and loudspeaker array
The attack range is a critical factor affecting whether the inaudible attacks can be presented in practical scenarios; a more extended attack range can ensure the attacker’s concealment. Different victim devices can also affect the performance of MetaAttack, as the microphones and voice systems used by those have different recognition abilities. Therefore, to test the maximum attack range on various devices of MetaAttack, we let our system attack the devices (devices’ parameters are shown in Table III) from various ranges, and the experimental results are shown in Fig. 14. The outcomes demonstrate that MetaAttack exhibited optimal performance on the Apple iPhone 14 Pro, achieving a maximum attack range of 9.2 m (Success Rate = 90%). The finding suggests that MetaAttack excels particularly on Siri, potentially attributed to Siri’s superior denoising algorithm. Moreover, MetaAttack also delivered a strong performance on the other devices, with an average maximum attack range of 8.85 m (Success Rate greater than 50%). The analysis of Fig. 14 reveals that the coverage of Part A and Part B displayed in Fig. 2 is attained by all victim devices. Notably, the loudspeaker array achieved an average maximum attack range of only 2.72 m. The improved performance of MetaAttack can be attributed to the filtering and amplification capabilities of the MAM, which enables the loudspeaker array to operate at higher power and enhances its energy. These experimental results indicate that MetaAttack performs effectively across various victim devices, and its compact size further highlights its superior concealment and practicality.
Figure 14: Performance of MetaAttack on various devices, Part A and Part B refer to the quadrants in Fig 2
VI-B2 A2 - Attack range by different arrangements of loudspeaker arrays
Our simulations have shown that MetaAttack has the most effective beamforming when using 12 loudspeakers (see Section IV-C). However, experimental results indicate that the placement of the 12 loudspeakers affects the performance of MetaAttack. To identify the optimal arrangement for MetaAttack, six common settings (as shown in Fig. 17) were systematically tested in experiments, with the results shown in Fig. 17. We find that layout (a) facilitates the convergence and interference of the output sound waves from individual speaker units in a specific direction, generating a concentrated, narrow beam-like sound field. The directional nature of this sound field enhances the focusing of sound, mitigates scattering and reflection effects, and thus extends the propagation distance of the sound waves.
Figure 15: (a) Six different loudspeaker arrays. (b) Performance of loudspeaker arrays
Figure 16: Filtering performance of MetaAttack
Figure 17: Enhanced performance of MetaAttack
VI-B3 A3 - Filtering of audible leakage
To ensure the stealth of the attack, MetaAttack needs to achieve MetaAttack-side inaudibility for different inaudible commands. We use the as-k6 noise tester to test the filtering effects of MetaAttack at frequencies between 100 Hz - 4000 Hz. The experimental results are shown in Fig. 17. It has been observed that MetaAttack can effectively filter audible leakage both in front of, on sides of and behind it to levels below the “Leakage Threshold”. Moreover, the consistency between the experimental results and the filtering simulation results shown in Fig. 6 confirms the accuracy of our experimental results.
VI-B4 A4 - MetaAttack’s enhancement functionality
The enhancement functionality of MAM plays a pivotal role in enabling the reliance on 12 compact loudspeakers to exceed the long-range inaudible commands transmission achieved by large speaker array [14]. Through practical experimentation in real-world scenarios utilizing ablation techniques, we conducted comparative analyses on the impact of MAM and LAM (absent enhancement) on 12 distinct victim devices. The experimental outcomes, as shown in Fig. 17, demonstrate the substantial improvement facilitated by the enhancement feature of MAM, resulting in an average attack distance of 8.85 m, which represents a notable enhancement of 17.5% compared to LAM. Furthermore, an investigation was conducted to assess the efficacy of the inaudible commands transmitted by MAM. This was achieved through the use of the as-k6 noise tester. The results demonstrated that the energetic output of inaudible commands from MAM and LAM were 107 dB and 95 dB, respectively. These measurements serve as evidence of the superior performance capabilities of MAM.
VI-B5 A5 - Noise impact
Another crucial factor influencing the success rate of inaudible attacks is the ambient noise level. This factor is significant as noise can disrupt the accuracy of the words in the inaudible commands, leading to failure if crucial words are not correctly identified. To investigate this effect, we tested the attack range of MetaAttack across 12 distinct victim devices (consistent with A4) under varying noise conditions. The results of these experiments are illustrated in Fig. 19. In environments with noise levels approximately ranging from 55 to 65 dB (comparable to library or study areas), the maximum attack ranges of MetaAttack across different devices almost remain unaffected. However, as the noise level increases to approximately 70 to 75 dB (comparable to noisy office settings or urban traffic sounds), the maximum attack range of MetaAttack is reduced to approximately 7 - 8 m (a reduction of about 1 m). These findings suggest that the robust inaudible commands transmitted by MetaAttack demonstrate resilience against environmental noise, enabling successful inaudible attacks even in typical noise conditions.
VI-C Performance in Real-world Scenarios
Figure 18: The impact of different noise levels
Figure 19: The impact of different wind speeds
VI-C1 S1 - Long-range attacks between high-rise buildings
The propagation of ultrasound is influenced by the density of the medium, while changes in air density caused by wind speed impact the propagation of ultrasound. In contrast to the previous experiment A1, scenario S1 involved conducting covert and long-range attack experiments between two high-rise buildings, necessitating consideration of wind speed interference at high altitudes. In addition, the attacker is located in a public corridor with no stable power outlets nearby, and pedestrians frequently pass through the area, making it difficult to deploy large attack devices that require DC power. The efficacy of the MetaAttack in conducting attacks on target devices was evaluated under varying wind speed conditions in scenario S1, as depicted in Fig. 13a. As shown in Fig. 19, in typical wind speed environments (2 m/s - 3 m/s), the MetaAttack demonstrated a relatively stable performance, maintaining an average maximum attack distance of 8.6 m. When wind speeds exceeded 3.5 m/s, there was an impact on the attack distance. But attackers can achieve successful long-distance attacks by using weather forecasts or other means to avoid unfavorable wind conditions in advance.
VI-C2 S2 - Attacks against walking victims
Due to its portability, MetaAttack is well-suited for conducting covert attacks on walking victims in outdoor settings. In a real-world scenario illustrated in Fig. 13b, we recruited volunteers carrying various types of devices (specifically for barebones testing) for practical validation at different walking speeds. The volunteers with victimized devices maintained walking speeds of approximately 1 m/s, 1.5 m/s, corresponding to normal walking and brisk walking states, respectively. The experimental outcomes depicted in Fig. 20 demonstrate that when a victim moved at a speed of 1 m/s (average human walking pace) and was 5 m away from the MetaAttack, the success rate remained at 51.4%. It was observed that while MetaAttack maintains a consistent attack distance, the success rate of the attack significantly drops as the victim moves at a faster pace. Moreover, the efficacy of the MetaAttack varies across different victim devices. Related research [55, 56, 57] indicates that a distance of 3-4 m between individuals signifies a public distance suitable for interactions with unfamiliar individuals. Consequently, we believe that under regular walking circumstances, a trailing attack using MetaAttack beyond the three-meter threshold is unlikely to alert the victim, facilitating ease of success. To confirm the feasibility of MetaAttack, we released attack video against walking victims [32].
Figure 20: The impact of moving target’s speeds (target’s speeds are 1 m/s and 1.5 m/s)
VI-C3 S3 - Eavesdropping attacks in indoor environments
Eavesdropping is a prevalent method of information theft in commercial activities, involving covertly monitoring private conversations or communications to obtain sensitive or confidential information for unauthorized purposes. In S3 (as shown in Figure 13c), attackers can accomplish eavesdropping by sending inaudible commands to the victim’s phone to initiate a call to a number pre-set by the attacker. By leveraging the victim devices for eavesdropping, there is no necessity to physically retrieve the eavesdropping device from the premises, thereby mitigating the risk of exposure. Nonetheless, the primary challenge in this scenario lies in necessitating enhancements in character recognition accuracy. Therefore, we conducted tests to measure the word accuracy of MetaAttack in transmitting the command “call 1234567890” to various victims’ devices across different distances. As depicted in Fig. 24, the experiments indicated that MetaAttack achieved 100% word accuracy when targeting an Apple iPhone 14 Pro at a set attack distance of 8.7 m. For other devices, MetaAttack achieved 100% word accuracy within an average attack range of 8 m, demonstrating its capability to conduct eavesdropping attacks in most indoor settings.
VI-C4 S4 - Effect of passing individuals on attack range
Passing individuals can significantly impact the propagation of inaudible sound waves. To investigate the influence of obstacles on the range of MetaAttack attacks at various distances, we conducted experiments using a public study room setting as an illustrative scenario. The detailed experimental parameter settings are shown in Fig. 13d. The experiment involved two passing individuals. As shown in Fig. 23, taking the iPhone 14 Pro as an example, when someone crosses MetaAttack’s attack path back and forth at a speed of 1 m/s, the system’s performance is minimally affected, allowing the attack to reach average distances of up to 7.6 m. Experimental findings indicate that passing individuals absorb and scatter ultrasonic waves, gradually attenuating the transmitted waves. But it will not significantly affect the attack range of MetaAttack.
Figure 21: Effect on attack range by passing people
Figure 22: Attack Range vs. Number of Speakers
Figure 23: The feedback accuracy of MetaAttack for different scenarios and devices
VI-C5 S5 - Feedback on the results of an inaudible attack
The feedback mechanism is crucial for inaudible attacks, as it enables the attacker to swiftly clean the attack site to evade detection by the victim. In this experiment, we evaluated the efficacy of MetaAttack’s feedback across scenarios S1 to S4. The experimental results, depicted in Fig. 23, demonstrate varying Precision and Recall values across different devices. Notably, MetaAttack exhibited enhanced performance when victim devices were Apple. Conversely, MetaAttack exhibits poorer performance in unstable scenarios (Figures 13a and 13b), indicating that external factors have a greater impact on the feedback mechanism. In such instances, the attacker must exercise timely judgment based on personal experience to avoid detection by the victim, particularly in the event of a successful and continued attack.
Figure 24: The word accuracy of MetaAttack for different devices
VI-D Performance Compared to Prior Research
We made a comparison between the MetaAttack and existing primary attack devices. Theoretically, increasing the number of loudspeakers can potentially increase the attack range; however, too many loudspeakers can compromise concealment. In Fig. 23 we show the correlation between the attack range and the number of loudspeakers. An important improvement of the MetaAttack over previous studies is its invisibility and portability. As shown in Fig. 23, the MetaAttack uses only 12 speakers for its attacks and relies on only two 70 mm $\times$ 130 mm lithium batteries for power. LipRead [14] used 61 speakers for its operations and relied on power from a DC power supply. Another system called EchoAttack [9] has a limited attack range of 2.23m with 40 speakers. The enhanced DolphinAttack[15] has an extended attack range of 19.8 m, targeting the iPhone X. These enhancements are achieved through the use of 40 speakers and external accessories. However, this setup may affect the feasibility of performing covert attacks. These results show that MetaAttack is superior in terms of portability and execution of covert attacks, in contrast to previous research.
VII DISCUSSION
Hardware devices for inaudible attacks. Investigating the influence of distinct attack devices on the functionality of MetaAttack system represents an interesting direction for future research. Through experimentation with a range of speaker and amplifier models, it is expected that MetaAttack system may potentially enhance its attack range efficiency while optimizing resource consumption. This aspect is worthy of further exploration and analysis in future research endeavors.
Attack through walls, windows, or curtains. The attenuation of ultrasonic signals when passing through solid objects poses a significant challenge for transmitting inaudible commands at 40 kHz. This makes it difficult for MetaAttack to effectively attack victims in enclosed environments. Future work will focus on addressing this issue.
Possible conflicts in the feedback mechanism. Considering that a few attack commands (such as ”turn on airplane mode”) may conflict with feedback mechanism, we can first confirm the position of the victim device through feedback mechanism before executing the attack.
VIII COUNTERMEASURES
Previous defense strategies. The previous defense strategies against inaudible attacks can be mainly divided into two categories: forensics-based defense strategies and cancellation-based defense strategies. Forensics-based defense strategies utilize the nonlinear characteristics of inaudible commands to identify them by detecting specific differences between these commands and legitimate commands [8, 14]. In contrast, cancellation-based defense strategies by using additional ultrasonic speaker arrays to emit sound waves that cancel out the inaudible commands [58].
Challenges and Future Directions in Defending against MetaAttack Attacks. Forensics-based defense strategies can only issue warnings upon detecting an attack, while attacks from MetaAttack can often succeed within a few seconds, making it difficult for victims to timely stop ongoing attacks [58]. Although cancellation-based defense strategies can effectively prevent ongoing attacks, their effective range is limited to just a few meters. Thus, when defending MetaAttack that can conduct outdoor mobile attacks, victims must carry cumbersome equipment, such as speaker arrays and laptops, which increases inconvenience. We believe that future research should focus on developing more portable protective devices that can be carried easily or directly installed on smart devices to address the threats posed by MetaAttack.
IX RELATED WORK
IX-A Attacks on Voice Control Systems
Attacks on voice control systems have been proven to cause serious security issues. The previous attacks are mainly into two types: attacks based on adversarial examples and inaudible attacks. The attacks based on adversarial examples are presented by deceiving the victims [59, 60, 61]. However, these attacks based on adversarial examples rely on audible sounds, which can cause suspicion among alert victims. Inaudible attacks can directly disrupt or control voice control systems through ultrasound, which makes them more covert [7, 8, 9, 14]. For example, DolphinAttack [8] modulates attack commands onto ultrasound waves for attack, achieving control over smart devices. LipRead [9] and DolphinAttack (Long-range System) [15] expands the range of inaudible attacks on this basis, but its large devices are easily detected. Our MetaAttack first combines attacks with acoustic metamaterials to achieve long-range attacks in a compact size, making it more covert compared to previous studies.
IX-B Acoustic Metamaterials
Acoustic metamaterial designed for ventilation barriers is an emerging research field. This acoustic metamaterial can achieve the filtering function while maintaining ventilation through compact size. For example, Ghaffarivardavagh et al. [42] proposed a deep subwavelength acoustic metamaterial using Fano-like resonance. Subsequently, Zhu et al. [25] increased the filtering bandwidth of the acoustic metamaterial through nonlocal effects (600 Hz - 1900 Hz). Based on the above research, MetaAttack further expands the filtering range of this type of metamaterials (100 Hz - 4000 Hz) and first expands its application field to the inaudible attacks.
X CONCLUSION
This paper has presented the first inaudible attack based on acoustical materials. This new attack significantly extends the attack range while maintaining a compact setup. We have enhanced conventional acoustical materials in two ways. Firstly, we optimized the structure of the acoustic metamaterial to filter out audible leakage resulting from inaudible attacks effectively. Secondly, we transformed the acoustic metamaterial into a piston sound source, enabling the use of a small speaker array to achieve the same beamforming effect as larger devices. Importantly, our methodology utilizes commercially available off-the-shelf components, ensuring practical implementation. We tested the performance of our design under various experiments against prior methods. Our approach can achieve covert and longer-range inaudible attacks using a smallest device compared to prior works.

### v2

@@
 We showcase that MetaAttack can be used to launch inaudible attacks for representative voice-controlled personal assistants, including Siri, Alexa, Google Assistant, XiaoAI, and Xiaoyi.
-The average word accuracy of all assistants is 76%, with a range of 8.85 m.
-Index Terms:
-Inaudible Attack, Acoustic Metamaterials, Voice Control Systems, Security and Privacy.
-publicationid: pubid: $0$–$00$ 2021 IEEE
+The average success rate of all assistants is 76%, with a range of 8.85 m.
+publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
 I INTRODUCTION
-Voice control is a natural way for human-computer interactions, commonly found in mobile and smart home applications [1, 2, 3, 4, 5, 6].
+Voice control is a natural way for human-computer interactions, commonly found in mobile and smart home applications [1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
 Given its capacity to control a broad spectrum of tasks – from making phone calls to managing devices like surveillance cameras and even initializing bank transfers – voice control systems have increasingly become the target of attacks [11, 12, 13, 14, 15, 16, 17, 1].
@@
 Such attacks modulate voice commands onto ultrasonic carriers, generating commands imperceptible to human ears but can be accepted by voice control systems.
-Despite some levels of success in specific scenarios, existing attacking methods have a short attack range of up to 3 m.
+Despite some levels of success in specific scenarios, existing portable inaudible attacking methods (such as relying on lithium batteries and using small speaker arrays to transmit attack signals) [13, 18, 12] have a short attack range of up to 3 m.
 However, a close-in attack within 3 m can be easily detected by many vigilant individuals [19].
 Extending this range by simply increasing transmission power would lead to audible noise, thereby compromising the stealthiness of the attack.
-More recent work, as represented by LipRead [14], divides the spectrum of the inaudible commands to reduce audible leakage and utilizes loudspeaker arrays to expand the attack range.
-LipRead is shown to be able to launch inaudible attacks from distances up to 7.62 m using 61 loudspeakers.
-Another representative system, DolphinAttack [15], can reach 19.8 m for one model of the device (iPhone X), requires 40 speakers, and uses an external power supply and amplification equipment.
-However, the division of the spectrum of inaudible commands reduces the accuracy of inaudible commands, and the size of the amplification equipment makes it difficult to launch the attack in many practical public scenarios.
+To address this limitation, more recent work, as represented by LipRead [19], divides the spectrum of inaudible commands to reduce audible leakage and employs a large array of 61 speakers to extend the attack distance up to 7.62 meters.
+DolphinAttack [20], targeting the iPhone X, uses 40 speakers along with external power supplies and amplifiers to achieve a range of 19.8 meters.
+In comparison, although these methods can significantly expand the attack distance, their reliance on bulky equipment and external power sources greatly reduces portability, limiting their applicability in many real-world public settings.
 This paper demonstrates the feasibility of executing long-range, inaudible attacks through a compact audio modulation system that can easily fit into a modest-sized travel bag.
@@
 1 depicts a possible attack scenario where MetaAttack is covertly placed in a public place next to a window to target a device located in another building within 8.5 m.
-Figure 1: A possible attacking scenario of MetaAttack where an attacker can launch an inaudible attack in a public place to remotely control the victim device in another building.
+Figure 1: A possible attacking scenario of MetaAttack where an attacker can launch an inaudible attack in a public place to remotely control the target device in another building.
 In this scenario, the attack device must be stealthy and does not require the use of other continuously powered devices such as DC power supply.
@@
 We showcased that MetaAttack can be implemented with low-cost, commercial, off-the-shelf components like Raspberry Pi, amplifiers, loudspeakers, servo, and a 3D-printed MAM.
-We thoroughly evaluate MetaAttack under different practical scenarios under different attack ranges, tasks, noise levels, obstacles, and targeting different victim devices.
+We thoroughly evaluate MetaAttack under different practical scenarios under different attack ranges, tasks, noise levels, obstacles, and targeting different target devices.
 We compare MetaAttack against three prior inaudible attacks [12, 13, 19].
-Experimental results show that MetaAttack is robust under different evaluation scenarios and tasks and can successfully launch an audible attack within a 8.85 m range with a noise level of up to 65 dB.
-Compared with prior attacking methods, MetaAttack achieves the good accuracy and attack range (76% average accuracy under a 8.85 m attack range) with the smallest sized device.
+Experimental results show that MetaAttack is robust under different evaluation scenarios and tasks and can successfully launch an inaudible attack beyond 8 m range with a noise level of up to 65 dB.
+Compared with prior attacking methods, MetaAttack achieves a 76% average success rate under a 8.85 m attack range, making it the smallest device to achieve such performance.
 This paper makes the following contributions:
@@
 The concept behind these attacks is to issue commands to voice-activated devices without the device owner’s knowledge.
-Specifically, when microphones used in voice control systems receive input signals exceeding 20 kHz, due to the non-linearity effect, the microphones output additional non-linear signals ${S_{in}}^{2}(t)$ [15].
-${S_{in}}(t)$ as shown in Eq.1:
+In an inaudible attack, the microphone receives input signals that include not only the audible-frequency voice control signal but also a high-frequency carrier signal.
+Due to the microphone’s non-linear response, these input signals generate additional non-linear components in the output, providing the conditions for recognizing and transmitting the attack signal.
+Eq.1 represents the composition of the input signal ${S_{in}}(t)$, which is as follows:
 ${S_{in}}(t)={V_{attack}}\left(t\right)\cos\left({2\pi{f_{carrier}}t}\right)+\cos\left({2\pi{f_{carrier}}t}\right)$ (1)
-where ${V_{attack}}\left(t\right)$ is the voice control signal, and $\cos\left({2\pi{f_{carrier}}t}\right)$ is the high frequency carrier signal.
-When the attack signal is input to the microphones, the microphones output the signal $S_{attack}(t)$ as shown in Eq.2:
+where ${V_{attack}}\left(t\right)$ is the voice control signal, typically a low-frequency control signal that the device can recognize.
+$\cos\left({2\pi{f_{carrier}}t}\right)$ is the high-frequency carrier signal, with a frequency of $f_{carrier}$}, which is beyond the human hearing range (typically above 20 kHz).
+The role of the carrier signal is to add additional frequency components to the attack signal, utilizing the non-linear characteristics of the speaker array when outputting the signal to generate additional components.
+The attack signal output by the speaker array, ${S_{attack}}(t)$, is represented by Eq.2:
 $S_{attack}(t)$ $=S_{h}+\frac{A_{2}}{2}\left(1+V_{attack}^{2}(t)+2V_{attack}(t)\right)$ (2)
-where ${V_{attack}}\left(t\right)$ can be recognized by the microphones, allowing the inaudible attacks to be successfully realized.
+where $S_{h}$ represents background noise or other unrelated signals that are not associated with the attack.
+The term $\frac{A_{2}}{2}\left(1+V_{attack}^{2}(t)+2V_{attack}(t)\right)$ represents the additional signal components introduced by non-linear effects.
+In this case, ${V_{attack}}\left(t\right)$ can be recognized by the target’s microphone, enabling the successful execution of the inaudible attack.
 By utilizing the modulation method described above, an adversary encodes commands in ultrasonic frequencies that the microphones of nearby voice-activated devices can pick up.
@@
 However, the voice recognition of the targeted device can process these sounds as normal voice commands and execute the action without the user’s knowledge or consent, posing a significant vulnerability for voice-controlled smart devices.
-The effectiveness of inaudible attacks is compromised by the inherent limitation of a limited attack range compared to laser [34] and wireless signal attacks [35], typically ten feet [8, 9].
-This limitation results from the audible leakage at the point of transmission during inaudible attacks.
-Concealment requires a reduction in output power [14], thereby limiting the range of inaudible attacks.
-As a result, the impact of an inaudible attack decreases significantly as the distance between the sonic emitter and the target increases, making it less suitable for long-range targeting than other forms of attack.
 II-B Threat Model
-We focus on implementing an inaudible acoustic attack that is portable, miniaturized, and out of detection range.
-These characteristics address the need for attacks in portable and stealthy environments.
-Our attack does not require the attacker to access the victim’s device physically.
-However, we assume the attacker can synthesize the target’s voice using, e.g., deep-learning-based voice generation from a few training samples [8].
-We believe this assumption is reasonable in many day-to-day scenarios where the victim’s voice samples can be collected from public places, a multi-user meeting, or audio and video clips available to the attacker.
-We also assume that the victim was not intact with the device during the attack and surroundings are not completely enclosed so that it is possible to transmit ultrasounds.
+Portable, miniaturized and stealthy inaudible attack.
+We focus on implementing an inaudible acoustic attack that is highly portable, miniaturized, and can be executed outside of detection range.
+Miniaturization and portability mean that the system does not rely on signal amplifiers powered by DC power supplies or bulky speaker arrays, and can be easily carried in a handbag or small backpack, allowing the attacker to operate discreetly in public spaces.
+The so-called detection range refers to a range in which the target is unlikely to detect the presence of the attacker.
+According to related studies [39, 40, 41], a distance greater than 3 meters is considered a safe distance when interacting with strangers, typically not raising suspicion.
+No victim device interaction during attack.
+We assume that during the attack, the victim’s device is in close proximity but not actively in use, with its voice assistant idle and the screen not being viewed by the victim.
+This assumption aligns with those made in prior audio-based attack studies, such as Dolphin [12, 20], NUIT [42] and GhostTalk [15], and reflects realistic scenarios, such as when the victim places the device aside or face-down on a desk to focus on tasks like studying or attending a meeting.
+Obtain the victim’s voiceprint.
+We assume the attacker can synthesize the target’s voice using, e.g., deep-learning-based voice generation from a few training samples [12].
+We believe this assumption is reasonable in many day-to-day scenarios where the target’s voice samples can be collected from public places, a multi-user meeting, or audio and video clips available to the attacker.
+Attack victim device in pocket.
+We consider the scenario where the victim may place their device in their pocket for convenience while out, during which the pocket could physically interfere with the transmission of attack commands.
+For this scenario, we have verified the system’s effectiveness in Section VI-C2.
 II-C Practical Limitations of Prior Methods
-Launching inaudible voice commands involves modulating the control voice signal.
-However, nonlinear effects from devices like loudspeakers can cause audible leakage, particularly when operating at high power levels, to extend the attack range.
-Previous works [8, 9] limited device power to prevent sound leakage, but doing so limits the attacking range to around 3 m.
-To enable longer-range inaudible attacks, LipRead [14] segments the inaudible command spectrum into low-frequency parts that are less detectable by humans.
-While representing an improvement, LipRead also has drawbacks.
-It requires an array of 61 loudspeakers to reach up to an attacking distance of 7.62 m, increasing the risk of detection, and the method of segmenting the spectrum may lower command accuracy due to potential recombination or interference during transmission.
-Similarly, although Yan et al.
-[15] were able to extend the ultrasonic attack range to 19.8 m, the attack needed a signal generator and a constant DC power supply, which makes it less practical for portable mobile attacks.
-Additionally, the use of 40 loudspeakers was necessary for the attack device.
-To more clearly illustrate the limitations of existing studies regarding the distance of the inaudible sound attack, we present a visual representation in Table I.
-Most of the previous research has been on defeating voiceprint recognition [14, 7, 9, 8], but these studies are difficult to conduct in real-world scenarios.
-Some of the work requires close proximity to the attacking device, and although long-range attacks can be achieved, they require the introduction of more auxiliary devices [15, 14, 7] that are not stealthy.
-A viable attack therefore requires a balance between lightness and range.
+Launching inaudible voice commands is achieved by modulating the control signal, but high-power operation of speakers can cause nonlinear effects, leading to audible leakage, which limits the attack range.
+Previous studies [12, 13] avoided leakage by limiting the device power, but this restricted the attack range to within 3 meters.
+To achieve longer-range attacks, LipRead [19] divides the frequency spectrum into low-frequency parts to reduce human perception, but it requires 61 speakers, increasing the risk of detection, and spectrum segmentation may reduce the accuracy of commands.
+Similarly, Yan et al.
+[20] extended the attack range to 19.8 meters, but this requires a signal generator and DC power supply, which does not meet the needs of portable attacks, and requires 40 speakers.
+To clearly illustrate the limitations of existing research, we provide a comparison in Table I.
+Most studies focus on voiceprint recognition attacks [19, 11, 13, 12], but these attacks often require close proximity or a large number of auxiliary devices, lacking stealth.
+Therefore, a successful attack requires a balance between portability and attack range.
 TABLE I: Comparison of prior inaudible attack research in terms of attack range, portability and stealthy
-System name System setup DC power supply Signal generator Speaker number Attack range Portability Stealthy
-DolphinAttack (Long-range System) [15] [Uncaptioned image] Yes (Approximately 11 MacBook Pro) Yes (Approximately 12 MacBook Pro) 40 19.8m (iPhoneX) 10m (iPhoneSE) No No
-LipRead [14] [Uncaptioned image] Yes Yes (Approximately 7 MacBook Pro) 61 7.62m No No
-BackDoor [7] [Uncaptioned image] Yes Yes (Approximately 7 MacBook Pro) 18 3.5m No No
-EchoAttack [9] [Uncaptioned image] No Yes (Approximately 7 MacBook Pro) 40 2.23m Yes Yes
+System Setup Power supply Signal generator Speakers Attack range Portable Stealthy
+DolphinAttack (Long-range System) [20] [Uncaptioned image] Yes ($Width*Hight*Depth\approx$ $26*16.5*30~cm\approx$ 12MacBookPro)12~MacBook~Pro) Yes ($Width*Hight*Depth\approx$ $42.6*8.8*48.3~cm\approx$ 16MacBookPro)16~MacBook~Pro) 40 19.8m (iPhoneX) 10m (iPhoneSE) No No
+LipRead [19] [Uncaptioned image] Yes ($Width*Hight*Depth\approx$ $1.3*31.2*36~cm\approx 22$MacBookPro)22~MacBook~Pro) Yes ($Width*Hight*Depth\approx$ $26.1*10.3*30.3~cm\approx$ 7MacBookPro)7~MacBook~Pro) 61 7.62m No No
+BackDoor [11] [Uncaptioned image] Yes ($Width*Hight*Depth\approx$ $1.3*31.2*36~cm\approx 22$MacBookPro)22~MacBook~Pro) Yes ($Width*Hight*Depth\approx$ $26.1*10.3*30.3~cm\approx$ 7MacBookPro)7~MacBook~Pro) 18 3.5m No No
+EchoAttack [13] [Uncaptioned image] No Yes ($Width*Hight*Depth\approx$ $26.1*10.3*30.3cm\approx$ 7MacBookPro)7~MacBook~Pro) 40 2.23m Yes Yes
 VRIFLE [18] [Uncaptioned image] No No 8 1.8m Yes Yes
@@
 As shown in Fig. 3, MetaAttack is a 3-step inaudible attack, described in detail below.
-Figure 3: Overview of MetaAttack with three steps: Processing inaudible commands, Attacking victim’s device, and Activating feedback mechanism.
+Figure 3: Overview of MetaAttack with three steps.
 Processing inaudible commands.
-The initial phase of our attack is to process and modulate voice commands.
-In this work, we first captured the target’s voice and synthesize the target’s voiceprint (see Section II-B).
-Then we modulated the voice command groups generated by Google Text-to-Speech AI [39] and voiceprints onto ultrasound carrier waves using the modulation method discussed in Section II to form inaudible commands [14].
-Each voice command group used for the attack contains three types of inaudible instructions: commands for muting the victim devices, commands for executing attacks, and commands for feedback.
-For instance, mute commands might include “speak 6% ($MuteCmd$)” for Siri or “turn the volume to 1 ($MuteCmd$)” for other devices [35, 40], attack commands might include “send message to person X ($AttackCmd$),” while feedback commands could be “turn on hotspot ($FeedbackCmd_{1}$)” and “turn off hotspot ($FeedbackCmd_{2}$)”.
-Subsequently, a group of inaudible commands is transmitted through a loudspeaker array.
-During this process, any audible leakage is filtered out, and our MAM amplifies the inaudible commands.
-Attacking victim’s device.
-MetaAttack uses beamforming to send the inaudible commands, but this requires the signal to be transmitted within a specific angle of approximately 12°.
-To position the signal transmission angle, we use a servo to drive MetaAttack to point to the victim device within a 180° range in steps of 12°.
-We first send $MuteCmd$ to lower the victim device’s loudspeakers to an inaudible level[35, 40].
-Then, we send $AttackCmd$ to achieve the purpose of the attack.
-Finally, we send $FeedbackCmd_{1}$ and $FeedbackCmd_{2}$ to activate the feedback mechanism.
+The initial phase of our attack involves processing and modulating voice commands.
+We first capture the victim’s voice and synthesize their voiceprint (see Section II-B).
+Then, we modulate the voice commands generated by Google Text-to-Speech AI [45] together with the synthesized voiceprint onto ultrasonic carrier waves using the method described in Section II, generating inaudible commands [19].
+Each command group includes four types of inaudible instructions: Mute commands (e.g., “speak 6%” for Siri or “turn the volume to 1” for other devices [42, 46], denoted as $MuteCmd$), Attack commands ($AttackCmd$), Feedback commands (e.g., “turn on hotspot” and “turn off hotspot,” denoted as $FeedbackCmd_{1}$ and $FeedbackCmd_{2}$) and Reset commands (e.g., “speak 50%” for Siri or “turn the volume to 5” for other devices, denoted as $ResetCmd$).
+Attacking target device.
+To accurately target the device, MetaAttack uses a servo motor to rotate in 12° steps within a specified range, gradually aligning with the victim.
+The system first sends the $MuteCmd$ to mute the device [42, 46], then sends the $AttackCmd$ to carry out the attack, activates the feedback mechanism to confirm the success of the attack, and finally uses the $ResetCmd$ to restore the device’s volume.
+Leveraging recent research findings[47, 48, 49], our system has successfully implemented the covert deletion of call and text message records, thereby significantly enhancing the attack’s stealth capabilities.
+To ensure the reliability of the attack execution, each command is sent consecutively five times.
+Since the execution time of each command is only a few seconds, repeated transmissions do not introduce any significant negative effects.
+Leveraging MetaAttack’s high single-command success rate of 76% at a distance of 8.85 meters, this strategy raises the theoretical final success rate of each command to 99.76%.
+In addition, we conducted 50 rounds of complete feedback processes to validate the execution success rate of each attack stage.
+Among them, the $MuteCmd$, AttackCmdAttackCmd, FeedbackCmdsFeedbackCmds and $ResetCmd$ each achieved a 98% success rate, with only one failure per command.
+The overall success rate of the full attack chain was 94%, with only three failures, demonstrating the high efficiency and stability of the attack process.
 Activating feedback mechanism.
-The feedback mechanism uses the Raspberry Pi to detect the state changes caused by the wireless signal switching command received by the target device, thereby evaluating whether the attack is successful.
-If the hotspot information of the device completes state transitions between appearance and disappearance in the detection list of MetaAttack after the attacker sends a command group, we consider the execution of the command group successful.
-Considering that the SSID list within the attack range is not static, the SSID information from weak wireless signals of non-target devices may randomly appear or disappear, we use iwlist [41] to monitor the SSID and signal strength of wireless signals in real-time.
-Specifically, when the SSID information from weak wireless signals appears or disappears, its signal strength typically shows a gradual increase or decrease.
-In contrast, the wireless signal from the victim device triggered by $FeedbackCmd$ does not undergo such gradual changes when it appears or disappears.
-Therefore, after sending $FeedbackCmd$, we use iwlist to identify signals that appear or disappear directly, thereby eliminating the interference from weak wireless signals in our assessment.
-As shown in Algorithm 1 and Fig.
-4, we determine the attack’s success by identifying the hotspots of smartphones or the Bluetooth connectivity of smart home devices.
-To illustrate the feedback mechanism for smartphones, we break it down into four sequential steps:
+This feedback mechanism uses a Raspberry Pi to monitor the state changes of the victim device after receiving a wireless switching command to determine whether the attack was successful.
+If the victim device’s hotspot signal undergoes a state transition of appearancedisappearancereappearance, the command execution is considered successful.
+The reason for using the sequence appearancedisappearancereappearance is that relying solely on appearancedisappearance may lead to false positives.
+If a non-victim device happens to activate its hotspot between the two feedback commands, the system may incorrectly interpret this as an attack success and prematurely terminate the attack process.
+To address this, we introduce appearancedisappearancereappearance to confirm the continuity of the response, as it is highly unlikely for a non-victim device to switch its hotspot twice within a short period.
+To further reduce false positives, we incorporate iwlist [50] to monitor hotspot SSIDs and signal strength in real time, addressing the random appearance or disappearance of weak signals from non-victim devices.
+Typically, such weak signals show gradual changes in strength when appearing or disappearing, whereas the signal from the victim device triggered by $FeedbackCmd$ exhibits sudden changes.
+Based on this difference, we filter out SSIDs with abrupt signal strength transitions, effectively eliminating interference and improving detection accuracy.
+As shown in Algorithm 1 and Fig. 4, to illustrate the feedback mechanism for smartphones, we break it down into four sequential steps:
 Step 1.
-After completing the transmission of the $MuteCmd$ and $AttackCmd$ (line 1), the Raspberry Pi uses iwlist to record the identified hotspot devices and stores the information in the L1 list (line 1).
-The system then proceeds to send the $FeedbackCmd_{1}$ to store the records in the L2 list (line  1).
+After transmitting the $MuteCmd$ and $AttackCmd$ (line 1), the Raspberry Pi uses iwlist to record the identified hotspot devices and stores them in the L1 list (line 1).
+Then, it sends the $FeedbackCmd$ and stores the new records in the L2 list (line 1).
+To prevent the system from repeatedly executing $MuteCmd$ and $AttackCmd$ while waiting for the success of $FeedbackCmd$, we introduced the concept of a time window.
+Considering that the system needs a certain amount of time to receive the victim device’s hotspot signal via the Raspberry Pi, no other commands will be sent within a time window after sending $FeedbackCmd$, thereby avoiding the repeated execution of commands.)
 Step 2.
 The system continues to send $FeedbackCmd_{2}$ and stores the received hotspot information in list L3 (line 1).
-If there are discrepancies in the hotspot signals ($dif_{1}$) between the L2 and L3 lists, the process moves to Step 3 (lines  11).
+If there are discrepancies in the hotspot signals ($dif_{1}$) between the L2 and L3 lists, the process moves to Step 3 (line 1).
 Otherwise, the MetaAttack is rotated to the next angle (line 1).
@@
 MetaAttack will sequentially send $FeedbackCmd_{1}$ and then instruct the Raspberry Pi to retrieve the L4 hotspot list (line  1).
-If there are differences in the hotspot signals ($dif_{2}$) between the L3 and L4 lists, and the intersection of $dif_{1}$ and $dif_{2}$ contains only one device, a successful stealth attack is detected (lines  11).
-If not, the MetaAttack will be adjusted to the next angle (line  1).
-Step 4. To maintain the stealth of the attack, this step focuses on restoring the victim to its original state before the attack (lines  11).
-After these four steps, the system’s attack feedback mechanism ensures the success rate of the inaudible attack.
+If there are differences in the hotspot signals ($dif_{2}$) between the L3 and L4 lists, a successful stealth attack is detected (lines 1 - 1).
+If not, the MetaAttack will be adjusted to the next angle.
+Step 4. To maintain the stealth of the attack, this step focuses on restoring the victim device to its original state before the attack (line 1).
+Optimization of feedback strategies in different scenarios.
+Considering the time overhead of the feedback algorithm and the limited angular coverage of a single attack, we have designed additional strategies to improve attack efficiency: for attacks on visible targets, the attacker can manually align the system with the target, completing the attack within 1 minute.
+In addition, for commands that provide direct feedback to the attacker (such as making a call for eavesdropping), we immediately confirm whether the call was successfully connected after the command is executed.
+If the connection is successful, the attack is stopped via SSH connection to the Raspberry Pi; if it fails, the subsequent feedback is halted and the system moves to the next angle to reduce time overhead.
+By tailoring strategies to different environments, the system achieves efficient and flexible attack execution.
 To confirm the feasibility of our attack process, we have released a complete process demonstration [36] and open-sourced the feedback code [38].
-Figure 4: Feedback overview of MetaAttack (After using $MuteCmd$ and $AttackCmd$ to achieve the purpose of inaudible attacks, using $FeedbackCmd_{1}$ and $FeedbackCmd_{2}$ to provide feedback)
-Input: $AttackCmd,\ FeedbackCmd_{1},\ FeedbackCmd_{2}$
-[…]
-[…]
-2       $Send\ MuteCmd\ and\ AttackCmd$
-3       $Activate\ iwlist$
+Figure 4: MetaAttack Overview: MuteCmdMuteCmd and AttackCmdAttackCmd initiate the attack, FeedbackCmdFeedbackCmd confirms success and ResetCmdResetCmd restores the target.
+1
+2foreach angleangle do
+3
+4 SendMuteCmd,AttackCmd,activateiwlistSend\ MuteCmd,\ AttackCmd,\ activate\ iwlist
+5
 6 feedbackMechanism()
-5       if $\textnormal{{feedbackMechanism(}}\textnormal{\emph{}}\textnormal{{)}}\ \textbf{==}\ true$  then
+7  if $\textnormal{{feedbackMechanism(}}\textnormal{\emph{}}\textnormal{{)}}\ \textbf{==}\ true\$then
 8     break
 9  end if
-8
+10
 11 end foreach
-10
-11Function feedbackMechanism():
-       // $Achieve\ feedback\ mechanism$
+12
+13
+14Function feedbackMechanism():
+15
 16 Step1()
-13       Step2($L_{1},L_{2}$)
+17
 18 return  Step2($L_{1},L_{2}$)
-15
-16
-17
-18Function Step1():
+19
+20
+21
+22Function Step1():
   // $Attempt\ to\ turn\ on\ victim^{\prime}s\ hotspot$
 23  $Receive\ hotspot\ list\ L_{1}$
-20       $Send\ FeedbackCmd_{1}\ and\ receive\ hotspot\ list\ L_{2}$
+24  $Send\ FeedbackCmd_{1}$
+25  $Wait\ time\ window\ and\ receive\ hotspot\ list\ L_{2}$
 26  $\textbf{return}\ L_{1},L_{2}$
-22
-23Function Step2($L_{1},L_{2}$):
-       // $If\ any\ hotspot\ signal\ disappears,\ moves\ to\ Step3$
-24       $Send\ FeedbackCmd_{2}\ and\ receive\ hotspot\ list\ L_{3}$
+27
+28Function Step2($L_{1},L_{2}$):
+  // Ifanyhotspotsignaldisappears,movestoStep3If\ any\ hotspot\ signal\ disappears,\ moves\ to\ Step3
+29  $Send\ FeedbackCmd_{2},\ receive\ hotspot\ list\ L_{3}$
 30  $dif_{1}=L_{2}-L_{3}$
-26       if $dif_{1}.length\neq 0$ then
-27             Step3($L_{1},L_{3},dif_{1}$)
+31  if $L_{2}\neq L_{3}$ then
 32     return  Step3($L_{1},L_{3},dif_{1}$)
@@
 35  end if
-32
-33
-34
-35Function Step3($L_{1},L_{3},dif_{1}$):
-       // $If\ hotspot\ signal\ reappears,\ attack\ is\ successful$
-36       $Send\ FeedbackCmd_{1}\ and\ receive\ hotspot\ list\ L_{4}$
+36
+37
+38
+39Function Step3($L_{1},L_{3},dif_{1}$):
+  // Ifhotspotsignalreappears,attackissuccessfulIf\ hotspot\ signal\ reappears,\ attack\ is\ successful
+40  $Send\ FeedbackCmd_{1}$
+41  $Wait\ time\ window\ and\ receive\ hotspot\ list\ L_{4}$
 42  $dif_{2}=L_{4}-L_{3}$
 43  $Target\ list=dif_{1}\cap dif_{2}$
-39       if $Target\ list.length==1$ then
+44
 45 Step4($Target\ list,L_{4},L_{1}$)
-41             $\textbf{return}\ true$
-42      else
-43             $\textbf{return}\ false$
-44
+46  return true
+47
+48
+49Function Step4($Target\ list,L_{4},L_{1}$):
+  // RestorethehotspotandvolumestatusRestore\ the\ hotspot\ and\ volume\ status
+50  if $L_{4}-L_{1}==Target\ list$ then
+51     $Send\ FeedbackCmd_{2}\ and\ ResetCmd$
+52     ClearcallandtextmessagerecordsClear\ call\ and\ text\ message\ records
+53
 54  end if
-46
-47
-48Function Step4($Target\ list,L_{4},L_{1}$):
-       // $Restore\ the\ hotspot\ status\ of\ the\ victim\ device$
-49       if $L_{4}-L_{1}==Target\ list$ then
-50             $Send\ FeedbackCmd_{2}$
-51
-52       end if
-53
-54
-Algorithm 1 The framework of attack overview
+55
+56
+Algorithm 1 Attack overview
 IV Metamaterial Design for Covert Long-Range Attack
 IV-A Design Goals and Challenges
-To implement a covert long-range inaudible attack system, we need to achieve two goals:
-Design Goal 1.
-Filter out audible leakage.
-Maintaining the stealth nature of the attack requires filtering out audible leakage in the 100 Hz - 4000 Hz frequency range since the energy of audible leakage is concentrated in this spectrum.
-This filtering process is critical to maintaining the stealth nature of inaudible attacks and preventing their detection by acoustic leakage.
-Design Goal 2.
-Increase the range of inaudible commands.
-To further achieve long-range attacks, we need to increase the beamforming effect of the inaudible acoustic attack command.
-This is because when beamforming is improved, the sound waves can be more precisely focused in the direction of a specific victim device, reducing energy dispersion and loss, which allows the sound waves to travel farther.
-To achieve these two design goals, the system design must address two challenging issues.
+Design goal 1. Eliminate audible leakage in the 100 Hz - 4000 Hz range to maintain the stealthiness of the attack and avoid detection by acoustic leakage.
+Design goal 2. Enhance beamforming to focus inaudible commands more precisely on the target device, enabling longer-range attacks.
+To achieve these goals, the system must tackle two key challenges.
 Design challenge 1.
-How to design an acoustic metamaterial with a fixed geometrical structure, which ensures that the sound signal is not leaked (for Design Goal 1) but also allows the inaudible sound signal to propagate over a longer distance (for Design Goal 2).
+How to design an acoustic metamaterial, which ensures that the sound signal is not leaked (for Design Goal 1) but also allows the inaudible sound signal to propagate over a longer distance (for Design Goal 2).
 Design challenge 2.
@@
 IV-B Filter Out Audible Leakage
-To address Challenge 1, we introduce a novel Leak-shielding Acoustic Metamaterial (LAM).
+To address Challenge 1, we propose a novel Leak-shielding Acoustic Metamaterial (LAM).
 The basic filtering mechanism of the LAM is shown in Fig.
@@
 One question that naturally arises is how to determine the opening geometric size.
-Through our simulation in COMSOL, we find that the height (h) and the angle ($\gamma$) of the opening both affect the filtering range of LAM.
-Therefore, we have tried various combinations of h and $\gamma$, then as shown in Fig.
-5b, we find the opening height $h=3~{}mm$ and angle $\gamma={68.7^{\circ}}$ can expand the filtering range to 100 Hz - 4000 Hz.
-Figure 5: (a) The filtering principle of LAM. (b) LAM- Leak-shielding acoustic metamaterial
-Even though our LAM can utilize its unique structure to filter out audible leakage in front of the loudspeaker array through phase opposite (as shown in Fig.
-5a), it cannot prevent audible leakage from leaking from the sides or back of the loudspeaker array.
-To address this issue, our LAM has undergone another development:
+Through our simulations in COMSOL, we found that both the height (h) and the angle ($\gamma$) of the opening affect the filtering range of the LAM, with some key data from the simulation results shown in Fig.
+6.
+The leakage threshold in Fig.
+6 represents the sound pressure level (SPL) of audible leakage at different frequencies that can be heard by humans.
+According to previous studies, such as LipRead [19], the leakage threshold should be 5 dB below the threshold of hearing curve.
+Therefore, we derive the leakage threshold by subtracting 5 dB from the threshold of hearing curve [19].
+As shown in Fig.
+6a, the opening height (h) significantly affects the filtering performance in the low-frequency range.
+When $h=3$ mm, the sound pressure level (SPL) can be effectively suppressed below the leakage threshold.
+We believe this is because low-frequency sound waves, with their longer wavelengths, tend to propagate through the large-scale internal channels of the structure.
+In this case, the coupling strength between the central cylindrical structure (determined by h) and the rest of the system becomes critical (see Fig.
+5b).
+If h is too small, the coupling strength is insufficient, leading to poor filtering performance.
+If h is too large, excessive coupling may occur, resulting in excellent filtering performance at certain frequencies, while the filtering effect at other frequencies fails to effectively suppress leakage.
+A height of 3 mm strikes a good balance between reflection strength and frequency coverage, ensuring stable low-frequency filtering.
+In addition, as shown in Fig.
+6b, the aperture angle $\gamma$ is similarly crucial for high-frequency filtering.
+When $\gamma=68.7^{\circ}$, the SPL in this range can also be effectively suppressed below the leakage threshold.
+High-frequency sound waves, with shorter wavelengths, are more prone to strong multipath interference within the spiral blade structure.
+The angle $\gamma$ directly influences the coupling strength between this structure and the rest of the system (see Fig.
+5b).
+Similar to h, a $\gamma$ value that is too small or too large will cause filtering imbalance, whereas $\gamma=68.7^{\circ}$ achieves stable filtering in the high-frequency range.
+Therefore, we find the opening height h=3mmh=3~$m$ and angle $\gamma={68.7^{\circ}}$ can expand the filtering range to 100 Hz - 4000 Hz.
+Figure 5: (a) The filtering principle of LAM. (b) LAM- Leak-shielding acoustic metamaterial.
+Although the LAM can filter out audible leakage in front of the loudspeaker, it still struggles to block leakage from the sides or back. To address this, we further optimized its structure:
 Preventing the audible leakage from leaking to other directions.
 As shown in Fig.
-5b.
-Firstly, to prevent audible leakage from leaking from the sides of the loudspeaker array, we increase the thickness (${L_{1}}$) and the length (${L_{3}}$) of the wall on both sides of the LAM ($H=50mm$) to wrap around the loudspeaker array.
-This causes audible leakage on both sides to be reflected by the wall.
-Secondly, to eliminate the audible leakage on the back, we have added a cover plate at the rear of the LAM and filled the front of the cover plate with melamine sponge (${L_{2}}$) so that the audible leakage is eliminated by the thermal viscosity loss caused by the melamine sponge.
-Where length ${L_{3}}=17.5~{}mm$ and its value equals the thickness of the loudspeaker array.
-To investigate the effects of different wall thicknesses $L_{1}$ and melamine sponge thickness $L_{2}$ on eliminating sides and back audible leakage, we simulated them in COMSOL.
-Since the vast majority of voice commands have the energy to 100 Hz - 4000 Hz, we set the frequency in the range of these commands.
-Then, we let the sound waves emitted by the sound source pass through the LAM and calculated the energy of the sound waves on the sides and back of the LAM, measured in sound pressure level (SPL).
-The simulation validation results of some representative $L_{1}$ and $L_{2}$ are shown in Figure 6a and Figure 6b.
-To ensure that the audible leakage remains inaudible to people, prior works, such as LipRead [14], have demonstrated that the “Leakage Threshold”, should be 5 dB below the “Threshold of Hearing Curve”.
-As a result, we derive the “Leakage Threshold” by subtracting 5 dB from the “Threshold of Hearing Curve”, as depicted by the red line in Figure 6.
-The results demonstrate that when the wall thickness $L_{1}$ is greater than 2 mm and the melamine sponge thickness $L_{2}$ is greater than 40 mm, audible leakage is below the “Leakage Threshold”.
-Therefore, on the premise of considering the filtration efficiency and compactness of LAM, we have determined the wall thickness ${L_{1}}=2~{}mm$ and the melamine sponge thickness ${L_{2}}=40~{}mm$.
-The total length of the LAM is ${L_{4}=H+L_{1}+L_{2}+L_{3}}=109.5~{}mm$.
-Figure 6: (a) Filtering effect on sides of LAM. (b) Filtering effect back of LAM. (c) Filtering effect in front of LAM.
-To verify the filtering effect of the LAM, we tested its filtering performance in COMSOL.
-The simulation validation results are shown in Figure 6c.
-The result demonstrates that the audible leakage after passing through the LAM filter is below the “Leakage Threshold”, indicating the effectiveness of our LAM in mitigating audible leakage.
+5b, first, to prevent audible sound waves from leaking out from the sides of the speaker array, we increase the thickness ($L_{1}$}) and length ($L_{3}$}) of the walls on the sides of the LAM to encase the speaker array.
+This ensures that the audible leakage from the sides is reflected back by the walls.
+Second, to eliminate the audible leakage reflected to the back, we fill the back of the LAM with melamine sponge ($L_{2}$}), which absorbs the sound waves through its heat-viscous loss.
+Notably, we first confirm that the length ${L_{3}}=17.5~mm$, which is equal to the length of the speaker array.
+To investigate the effects of different wall thicknesses $L_{1}$ and melamine sponge thickness $L_{2}$ on eliminating sides and back audible leakage, we performed simulations in COMSOL.
+The simulation results for some representative $L_{1}$ and $L_{2}$ in suppressing side and back leakage are shown in Fig.
+7a and Fig.
+7b.
+The results in Fig.
+7a and Fig.
+7b show that when the wall thickness $L_{1}>2~mm$ and the melamine sponge thickness $L_{2}>40~mm$, the audible leakage is below the leakage threshold.
+To ensure inaudibility, it is crucial to choose an appropriate thickness to enhance the compactness of the LAM.
+Therefore, we determined the wall thickness ${L_{1}}=2~mm$ and melamine sponge thickness ${L_{2}}=40~mm$, which balance performance and volume, with the total length of the LAM being ${L_{4}=H+L_{1}+L_{2}+L_{3}}=109.5~mm$.
+Figure 6: (a) Impact of opening height $h$ on LAM front filtering. (b) Impact of opening angle $\gamma$ on LAM front filtering.
+Figure 7: (a) Impact of sponge thickness on LAM back filtering. (b) Impact of wall thickness on LAM sides filtering.
 Filtering audible leakage against inaudible attacks only addresses the stealth of the attack.
@@
 The piston sound source can enhance the high-frequency sound waves through acoustic resonance [34], allowing a small loudspeaker array to perform the same beamforming effect as a large loudspeaker array to enhance inaudible commands.
-Figure 7: Enhancement effect of inaudible commands in LAM
-Figure 8: MAM (Multi-functional Acoustic Metamaterial ($H_{2}=127.5~{}mm$).
-As shown in Figure 8, LAM consists mainly of a barrier and a circular piston in the center, and its structure is similar to the piston sound source [29, 30].
+Figure 8: Enhancement effect of inaudible commands in LAM.
+Figure 9: Multi-functional Acoustic Metamaterial (MAM).
+As shown in Fig.
+9, LAM consists mainly of a barrier and a circular piston in the center, and its structure is similar to the piston sound source [34, 35].
 When the inaudible commands pass through the LAM, the circular barrier causes them to converge into the circular hole-like structure, thus enhancing them.
 However, from the simulation results, we found that the surrounding spiral path inside LAM would disrupt the incident plane wave of inaudible commands, failing to form a piston sound source (the prerequisite for transforming the LAM to a piston sound source is that the incident waveform of inaudible commands is a plane wave [55, 56]).
-TABLE II: The impact of various factors on the enhancement effect of inaudible commands
+TABLE II: The impact of various factors on the enhancement effect
 Number of Loudspeakers Range (R) Carrier Frequency Maximum SPL
@@
 16 18 mm 40200 Hz 142 dB
-Figure 9: The sound field of loudspeaker array
-Figure 10: The results of enhancement effect.
-Figure 11: The prototype of MetaAttack
-Figure 12: The covert MetaAttack
+Figure 10: Speaker array’s sound field
+Figure 11: SPL of enhancement.
+Figure 12: Prototype of system
+Figure 13: Covert system
 To solve this issue, we let the inaudible commands pass through a LAM under different conditions and observed its waveform.
@@
 Table II provides the optimal combinations of range (R) and carrier frequency under different numbers of loudspeakers.
-The highest gain is achieved when 12 loudspeakers are incident at 40200 Hz, positioned at a range of 18 mm (i.e., $R=18~{}mm$) from the circular hole-like structure.
+The highest gain is achieved when 12 loudspeakers are incident at 40200 Hz, positioned at a range of 18 mm (i.e., R=18mmR=18~$m$) from the circular hole-like structure.
 We also found that when using 14 loudspeakers or more, the SPL no longer increases.
@@
 The final structure is depicted in Fig.
-8.
+9.
 This final structure is referred to as the Multi-functional Acoustic Metamaterial (MAM).
@@
 The result is shown in Fig.
-12.
+13.
 We have proven that under this structure, the waveform of the loudspeakers is a plane wave when it reaches the circular hole-like structure.
@@
 As shown in Fig.
-12, we find that when inaudible commands meet the above conditions and pass through the MAM, the SPL increases by 11 dB, indicating that the energy of inaudible commands has been enhanced.
+13, we find that when inaudible commands meet the above conditions and pass through the MAM, the SPL increases by 11 dB, indicating that the energy of inaudible commands has been enhanced.
 V IMPLEMENTATION
 V-A MetaAttack Components
-The prototype of MetaAttack uses commercial devices and a 3D printed MAM.
-Fig.
-12 shows our prototype system.
-This metamaterial serves as the basis for MetaAttack, which filters out low-frequency audible sounds and amplifies inaudible commands.
-The Raspberry Pi acts as a terminal for MetaAttack, facilitating the transmission of commands to attack victims and control the servo.
-It also detects the hotspot or Bluetooth signal from the victim device to determine the attack’s success.
-The amplifier amplifies the ultrasonic inaudible commands transmitted by Raspberry Pi.
-The servo helps MetaAttack to attack over a long distance.
-A loudspeaker array (12 cm $\times$ 1.4 cm) in MAM transmits inaudible commands with a center frequency of 40 kHz.
+The prototype of MetaAttack uses commercial devices and a 3D-printed acoustic metamaterial (MAM), as shown in Fig.
+13.
+This metamaterial filters out low-frequency sounds and amplifies inaudible commands.
+The Raspberry Pi serves as the terminal, responsible for sending attack commands, controlling the servo motor, and running the feedback mechanism.
+The speaker array(12 cm × 1.4 cm) in the MAM transmits commands with a center frequency of 40 kHz.
 V-B Stealthiness and Portability
-The stealthiness and portability of inaudible attacks play a significant role in facilitating covert operations, evading detection, targeting diverse devices, and expanding the attack surface to maximize the effectiveness and reach of these malicious activities.
+The stealthiness and portability of inaudible attacks are crucial for covert operations and evading detection.
 As shown in Fig.
-12, the length, width, and height of the assembled MetaAttack are 12.75 cm, 11 cm, and 18.8 cm, respectively.
-Its compact size allows us to carry it conveniently.
-Therefore, to enhance the stealth capabilities of MetaAttack, we have consolidated all attack devices into a compact travel bag.
-The bag can be held in hand or slung over the shoulder when performing a silent attack.
-With its compact size and enhanced stealth features, the MetaAttack in its travel bag configuration is a powerful tool for silent and covert operations.
+13, the dimensions of the assembled MetaAttack are 12.75 cm × 11 cm × 18.8 cm, making it compact and easy to carry.
+To enhance its stealth capabilities, we have consolidated all the attack devices into a travel bag that can be held by hand or slung over the shoulder during silent attacks.
+This compact and discreet design makes MetaAttack in its travel bag configuration an effective tool for silent operations.
 VI EVALUATION
@@
 In our evaluation, we use twelve devices and the specifications of the devices are given in the Table III.
-TABLE III: Victim devices
+TABLE III: Target devices
 Manuf. Model OS Voice system
@@
 (d) S4
-Figure 13: Real attack scenarios: (a) S1: The attacker attacks the victim from another building at a range of 8.5 m, (b) S2: The attacker attacks the walking victim (the victim’s speed is 1 m/s and 1.5 m/s), (c) S3: The attacker sends the inaudible command: “Call 1234567890” to achieve eavesdropping purpose.
-(d) S4: The attacker attacks the victim in an obstacle environment.
+Figure 14: Real attack scenarios are as follows: (a) S1: The attacker launches a remote attack from a building 8.5 meters away; (b) S2: The attacker targets a walking victim outdoors; (c) S3: The attacker sends a call command to eavesdrop; (d) S4: The attacker operates in an environment with obstacles.
 Metrics.
@@
 We conducted a range of experiments, all of which, except experiment S3, used “Send a message to Jack” as the inaudible command to achieve the purpose of sending false information.
-First, experiments A1 to A5 were designed to evaluate the impact of different factors on MetaAttack.
-Second, experiments S1 to S4 were conducted to evaluate the attack performance and the accuracy of the feedback system in real-world scenarios.
-As can be seen from Fig.
-13, the suitable attack distances for the S1 - S4 experimental scenarios are all over 4 m, and all of them are situated in outdoor public areas or small private spaces, making them unsuitable for using large external devices.
-We believe that these are common scenarios, but it is difficult to fully adapt the existing methods, either the stealth requirement is met but the attack distance is not sufficient [7] [9] [36] [8], or the attacking device has a certain range but is not stealthy [14] [15].
-Finally, we compared MetaAttack to prior research on inaudible attacks to highlight its superiority.
-Detailed descriptions of the experiments can be found in Table V.
-Our experiments were approved and overseen by the University’s IRB(Institutional Review Board) to ensure compliance with data security measures and ethical guidelines for data processing.
+Experiments A1 to A5 evaluated the basic performance of MetaAttack and verified the contribution of the metamaterials.
+The basic performance tests include attack range (A1), selection of loudspeaker array arrangement (A2), and the impact of environmental noise (A3); the contribution of the metamaterials is assessed through their ability to suppress audible leakage (A4) and the performance improvements brought by MAM and LAM (A5).
+Experiments S1 to S5 evaluated the attack performance and the accuracy of the feedback system.
+As shown in Fig.
+14, the attack distances for scenarios S1 to S4 all exceed 4 meters and are located in outdoor or small private spaces, making them unsuitable for large external devices.
+Existing methods typically involve a trade-off between stealthiness and attack range [11, 13, 18, 12], whereas MetaAttack performs better in this regard.
+Detailed experimental settings can be found in Table V.
 TABLE IV: Evaluating metrics
 Metrics Description
-Success rate The ratio between the number of successful attacks and the total number of attacks (30 times per experiment) [10, 9].
-RSA The Range at which MetaAttack can Successfully Attack (with a success rate of over 50% ) [14].
-Word accuracy The ratio of the number of correctly recognized words in inaudible commands to all words [8, 14, 15].
-Precision The proportion of successful attacks correctly identified by the feedback mechanism to all successful attacks identified by the feedback mechanism [53, 54, 14].
-Recall The proportion of successful attacks correctly identified by the feedback mechanism to all truly successful attacks [53, 54, 14].
+Success rate Number of successful attacks / Total attempts (30 trials per device per round) [14, 13]. For Exp. A1(VI-B1).
+RSA The Range at which MetaAttack can Successfully Attack (with a success rate of over 50% ) [19]. For Exps. A2(VI-B2), A4(VI-B4), A5(VI-B5), S1(VI-C1), S2(VI-C2), S4(VI-C4).
+Word accuracy Proportion of correctly recognized words in inaudible commands [12, 19, 20]. For Exps. S3(VI-C3).
+Precision The proportion of successful attacks correctly identified to all successful attacks identified by the feedback mechanism [62, 63, 19]. For Exp. S5(VI-C5).
+Recall The proportion of successful attacks correctly identified to all truly successful attacks [62, 63, 19]. For Exp. S5(VI-C5).
 TABLE V: List of experimental setting
 Test objectives Label Test focus Description
-A1 (Sec.VI-B1) Range of attack on different victim devices between MetaAttack and loudspeaker array To test the impact of range and victim devices to MetaAttack, we let MetaAttack attack the various devices (mentioned in Table III) at different ranges.
-A2 (Sec.VI-B2) Attack range by different arrangements of loudspeaker arrays To find the optimal arrangement of the 12 speakers (see Sec.IV-C) for MetaAttack, we used six common settings of the speaker array to attack the victim devices.
-A3 (Sec.VI-B3) Filtering of audible leakage To test the practical filtering performance of MetaAttack in inaudible attacks, we tested the audible sound leakage in front of and back of MetaAttack at frequencies between 100 Hz - 4000 Hz.
-A4 (Sec.VI-B4) MetaAttack’s enhancement functionality To test the practical performance of MetaAttack’s enhancement functionality, we compared the performance of MAM, LAM and loudspeaker array.
-Performance under va- rious conditions A5 (Sec.VI-B5) Noise impact To test the impact of background noise on MetaAttack, we continuously increased the noise from 55 dB to 75 dB and tested MetaAttack’s performance in this environment.
-Performance in real- world scenarios S1 (Sec.VI-C1) Long-range attacks between high-rise buildings To test MetaAttack’s long-range attack performance in real-world attack scenarios, we attacked the victim devices inside another building, with the two buildings located 8.5 m apart (as shown in Fig.
-13a).
-S2 (Sec.VI-C2) Attacks against walking victims To test MetaAttack’s performance in attacking against walking victims, we attacked a victim who is moving on the street, and the victim’s speed is 1 m/s and 1.5 m/s (as shown in Fig.
-13b).
-S3 (Sec.VI-C3) Eavesdropping attacks in indoor environments To test MetaAttack’s eavesdropping performance, we let MetaAttack send “Call 123456780” to attack the victim devices to call the attacker’s number for eavesdropping (as shown in Fig.
-13c).
-S4 (Sec.VI-C4) Effect of passing individuals on attack range To test MetaAttack’s performance in public environments, we attacked the victim devices in a public room, and there are people along the attack path (as shown in Fig.
-13d).
-S5 (Sec.VI-C5) Feedback on the results of an inaudible attack To verify whether the feedback mechanism of MetaAttack is accurate, we ran it in S1 - S4 and calculated its Precision and Recall.
-Performance compared to prior research Sec.
-VI-D Advantages of the MetaAttack in terms of concealment To verify that MetaAttack can achieve more covert attacks, we compared the size of its devices with previous attack system.
+A1 (Sec.VI-B1) Attack range performance Attack range of MetaAttack was evaluated in comparison to a standalone speaker array.
+A2 (Sec.VI-B2) Differences in speakers’ layout Six common array layouts were tested to identify the optimal configuration (Sec. IV-C).
+A3 (Sec.VI-B3) Noise impact MetaAttack’s performance was tested by varying background noise from 55 to 75 dB.
+A4 (Sec.VI-B4) Leakage filtering performance The audible sound leakage of MetaAttack was measured across different directions from 100-4000 Hz, alongside tests on human perception.
+Basic performance evaluation A5 (Sec.VI-B5) Metamaterials’ enhancement Enhancement performance was evaluated by comparing MAM, LAM, and speaker array.
+Real-world performance evaluation S1 (Sec.VI-C1) Building-to-building attack Target is attacked from 8.5 m away in a building under varying wind speeds (Fig. 14a).
+S2 (Sec.VI-C2) Attack against outdoor targets The target walking outdoors was attacked, while also verifying the attack effectiveness with the device in a pocket (Fig. 14b).
+S3 (Sec.VI-C3) Eavesdropping attacks MetaAttack’s eavesdropping was assessed by triggering a call (Fig. 14c).
+S4 (Sec.VI-C4) Impact of passing individuals Attacks were conducted in a room with passing individuals to explore their impact (Fig. 14d).
+S5 (Sec.VI-C5) Feedback mechanism performance Measured the precision and recall of the feedback mechanism in S1-S4 scenarios.
+Compared to prior research Sec. VI-D MetaAttack’s concealment advantages MetaAttack’s device size and range was benchmarked against prior systems to assess its enhanced covert capabilities.
 VI-B Performance under Various Conditions
-VI-B1 A1 - Range of attack on different victim devices between MetaAttack and loudspeaker array
-The attack range is a critical factor affecting whether the inaudible attacks can be presented in practical scenarios; a more extended attack range can ensure the attacker’s concealment.
-Different victim devices can also affect the performance of MetaAttack, as the microphones and voice systems used by those have different recognition abilities.
-Therefore, to test the maximum attack range on various devices of MetaAttack, we let our system attack the devices (devices’ parameters are shown in Table III) from various ranges, and the experimental results are shown in Fig.
-14.
-The outcomes demonstrate that MetaAttack exhibited optimal performance on the Apple iPhone 14 Pro, achieving a maximum attack range of 9.2 m (Success Rate = 90%).
-The finding suggests that MetaAttack excels particularly on Siri, potentially attributed to Siri’s superior denoising algorithm.
-Moreover, MetaAttack also delivered a strong performance on the other devices, with an average maximum attack range of 8.85 m (Success Rate greater than 50%).
-The analysis of Fig.
-14 reveals that the coverage of Part A and Part B displayed in Fig.
-2 is attained by all victim devices.
-Notably, the loudspeaker array achieved an average maximum attack range of only 2.72 m.
-The improved performance of MetaAttack can be attributed to the filtering and amplification capabilities of the MAM, which enables the loudspeaker array to operate at higher power and enhances its energy.
-These experimental results indicate that MetaAttack performs effectively across various victim devices, and its compact size further highlights its superior concealment and practicality.
-Figure 14: Performance of MetaAttack on various devices, Part A and Part B refer to the quadrants in Fig 2
-VI-B2 A2 - Attack range by different arrangements of loudspeaker arrays
-Our simulations have shown that MetaAttack has the most effective beamforming when using 12 loudspeakers (see Section IV-C).
-However, experimental results indicate that the placement of the 12 loudspeakers affects the performance of MetaAttack.
-To identify the optimal arrangement for MetaAttack, six common settings (as shown in Fig.
-17) were systematically tested in experiments, with the results shown in Fig.
-17.
-We find that layout (a) facilitates the convergence and interference of the output sound waves from individual speaker units in a specific direction, generating a concentrated, narrow beam-like sound field.
-The directional nature of this sound field enhances the focusing of sound, mitigates scattering and reflection effects, and thus extends the propagation distance of the sound waves.
-Figure 15: (a) Six different loudspeaker arrays. (b) Performance of loudspeaker arrays
-Figure 16: Filtering performance of MetaAttack
-Figure 17: Enhanced performance of MetaAttack
-VI-B3 A3 - Filtering of audible leakage
-To ensure the stealth of the attack, MetaAttack needs to achieve MetaAttack-side inaudibility for different inaudible commands.
-We use the as-k6 noise tester to test the filtering effects of MetaAttack at frequencies between 100 Hz - 4000 Hz.
-The experimental results are shown in Fig.
-17.
-It has been observed that MetaAttack can effectively filter audible leakage both in front of, on sides of and behind it to levels below the “Leakage Threshold”.
-Moreover, the consistency between the experimental results and the filtering simulation results shown in Fig.
-6 confirms the accuracy of our experimental results.
-VI-B4 A4 - MetaAttack’s enhancement functionality
-The enhancement functionality of MAM plays a pivotal role in enabling the reliance on 12 compact loudspeakers to exceed the long-range inaudible commands transmission achieved by large speaker array [14].
-Through practical experimentation in real-world scenarios utilizing ablation techniques, we conducted comparative analyses on the impact of MAM and LAM (absent enhancement) on 12 distinct victim devices.
-The experimental outcomes, as shown in Fig.
-17, demonstrate the substantial improvement facilitated by the enhancement feature of MAM, resulting in an average attack distance of 8.85 m, which represents a notable enhancement of 17.5% compared to LAM.
-Furthermore, an investigation was conducted to assess the efficacy of the inaudible commands transmitted by MAM.
-This was achieved through the use of the as-k6 noise tester.
-The results demonstrated that the energetic output of inaudible commands from MAM and LAM were 107 dB and 95 dB, respectively.
-These measurements serve as evidence of the superior performance capabilities of MAM.
-VI-B5 A5 - Noise impact
-Another crucial factor influencing the success rate of inaudible attacks is the ambient noise level.
-This factor is significant as noise can disrupt the accuracy of the words in the inaudible commands, leading to failure if crucial words are not correctly identified.
-To investigate this effect, we tested the attack range of MetaAttack across 12 distinct victim devices (consistent with A4) under varying noise conditions.
-The results of these experiments are illustrated in Fig.
-19.
-In environments with noise levels approximately ranging from 55 to 65 dB (comparable to library or study areas), the maximum attack ranges of MetaAttack across different devices almost remain unaffected.
-However, as the noise level increases to approximately 70 to 75 dB (comparable to noisy office settings or urban traffic sounds), the maximum attack range of MetaAttack is reduced to approximately 7 - 8 m (a reduction of about 1 m).
-These findings suggest that the robust inaudible commands transmitted by MetaAttack demonstrate resilience against environmental noise, enabling successful inaudible attacks even in typical noise conditions.
+VI-B1 A1 - Attack range performance between MetaAttack and loudspeaker array
+Attack range is a key factor in evaluating the practicality of ultrasonic attacks—greater range offers better concealment. Device microphones and voice systems also influence MetaAttack’s performance.
+We tested the maximum attack range of MetaAttack across different devices (results shown in Fig.
+15).
+The iPhone 14 Pro performed best, achieving a 90% success rate at a distance of 9.2 meters.
+Other devices had an average range of 8.85 meters with success rates above 76%.
+MetaAttack successfully covered Regions A and B in Fig.
+2 on all devices, while a regular speaker array had an average range of only 2.72 meters.
+The performance improvement is due to the MAM’s filtering and amplification capabilities, which significantly enhance signal strength.
+Figure 15: Success rate(SR) performance of MetaAttack on various devices, Part A and Part B refer to the quadrants in Fig 2.
+VI-B2 A2 - Differences in speaker’s layout
+Our simulation results show that the beamforming effect of MetaAttack is most significant when using 12 loudspeakers (see Section IV-C).
+However, we found that the speaker layout significantly impacts MetaAttack’s performance.
+To determine the optimal speaker layout, we simulated the sound pressure fields of six common configurations in COMSOL and systematically evaluated their RSA and SPL performance, as shown in Fig.
+16.
+As illustrated in Fig.
+16a, only layout (a) successfully forms an effective plane wave, while the others fail to do so.
+We attribute this to the fact that in layout (a), all speakers are positioned within the central circular opening of the metamaterial, avoiding interference from the surrounding spiral structures.
+Additionally, the speakers are evenly spaced and symmetrically arranged, which promotes coherent wavefront formation and facilitates the generation of a desirable plane wave.
+In contrast, the other layouts either suffer from obstruction by the spiral structures or lack consistent spacing and symmetry, which hinders effective wavefront shaping.
+Fig.
+16b further confirms that layout (a) achieves the best performance in both SPL and RSA, leading us to select it as the final design for the system.
+Figure 16: (a) Six speaker arrays with their sound fields. (b) Performance of speaker arrays (tests include arrays’ RSA and SPL at Transmitter(TX).
+Figure 17: (a) Leakage SPL comparison at TX (top) and its relation to the leakage threshold (bottom). (b) Volunteers’ perception of the leakage.
+Figure 18: (a) SPL comparison of inaudible attacks with and without MAM. (b) RSA performance comparison.
+VI-B3 A3 - Noise impact
+Another factor affecting the success rate of inaudible attacks is environmental noise, which can interfere with the accuracy of command recognition.
+We tested MetaAttack’s RSA on 12 devices under varying noise conditions.
+The results, shown in Fig.
+20, indicate that only when the noise level increases to 70-75 dB (similar to noisy office or urban traffic sounds) does the attack range decrease to 7-8 meters (a reduction of about 1 meter).
+This suggests that the strong beamforming effect of MetaAttack allows it to successfully perform inaudible attacks even in typical noisy environments.
+VI-B4 A4 - Leakage filtering performance
+To ensure the stealth of the attack, MetaAttack needs to effectively filter out audible leakage.
+In a laboratory environment with a background noise level of 30 dB, we played single-tone signals ranging from 100 to 4000 Hz at 60 dB SPL and used the as-k6 noise tester to evaluate the filtering performance of MetaAttack.
+As shown in Fig.
+17a, MetaAttack significantly attenuates sound waves from the front, sides, and rear, reducing audible leakage to below the Leakage Threshold.
+To verify whether the sound filtered by MetaAttack is imperceptible to the human ear, we invited 42 volunteers to participate in the experiment on a voluntary basis.
+The volunteers were evenly divided into three age groups (14 volunteers per group): 8–12 years old (children, more sensitive to sound), 13–18 years old (teenagers, broader hearing range), and 19–30 years old (young adults, peak auditory resolution), with a balanced gender distribution within each group.
+The experiment was conducted in an environment with a background noise level of 30 dB.
+Volunteers stood at various distances and attempted to identify whether there was any audible leakage at four levels: inaudible, mild noise, indistinct and intelligible [19, 12, 20, 64].
+As shown in Fig.
+17b, the experimental results indicate that only within a short distance of 0.1 to 0.5 meters could a few volunteers with more sensitive hearing perceive faint or indistinct sounds.
+We speculate that this may be due to slight audible leakage caused by nonlinear acoustic effects during air transmission.
+However, our attack range far exceeds this distance (reaching up to 8.85 meters), and the background noise level in the experimental environment was only 30 dB, which is extremely quiet.
+In real-world environments, such weak audio leakage is typically masked by higher background noise.
+Therefore, we believe that MetaAttack demonstrates strong imperceptibility in practical attack scenarios.
+VI-B5 A5 - Contribution of metamaterials
+The enhancement functionality of MAM plays a key role in enabling 12 compact loudspeakers to surpass large speaker arrays in achieving long-range inaudible command transmission [19].
+Through experiments conducted in real-world scenarios using ablation techniques, we compared the RSA of MAM, LAM and speaker array without metamaterial on 12 distinct target devices, and analyzed the SPL at various distances.
+As shown in Fig.
+18a, the results show that MAM (8.85 m) improves the average attack distance by 14.2% compared to LAM (7.75 m) and by 225% compared to the speaker array (2.72 m).
+Compared to LAM (95 dB), MAM (107 dB) increases the output SPL by 12.6%, and by 28.9% compared to the speaker array (83 dB).
+These results highlight the two key functions of MAM—audible leakage suppression and energy amplification—as essential to enhancing speaker array performance.
 VI-C Performance in Real-world Scenarios
-Figure 18: The impact of different noise levels
-Figure 19: The impact of different wind speeds
+Figure 19: Impact of noise levels.
+Figure 20: Impact of wind speeds.
 VI-C1 S1 - Long-range attacks between high-rise buildings
-The propagation of ultrasound is influenced by the density of the medium, while changes in air density caused by wind speed impact the propagation of ultrasound.
-In contrast to the previous experiment A1, scenario S1 involved conducting covert and long-range attack experiments between two high-rise buildings, necessitating consideration of wind speed interference at high altitudes.
-In addition, the attacker is located in a public corridor with no stable power outlets nearby, and pedestrians frequently pass through the area, making it difficult to deploy large attack devices that require DC power.
-The efficacy of the MetaAttack in conducting attacks on target devices was evaluated under varying wind speed conditions in scenario S1, as depicted in Fig.
-13a.
 As shown in Fig.
-19, in typical wind speed environments (2 m/s - 3 m/s), the MetaAttack demonstrated a relatively stable performance, maintaining an average maximum attack distance of 8.6 m.
-When wind speeds exceeded 3.5 m/s, there was an impact on the attack distance.
-But attackers can achieve successful long-distance attacks by using weather forecasts or other means to avoid unfavorable wind conditions in advance.
-VI-C2 S2 - Attacks against walking victims
-Due to its portability, MetaAttack is well-suited for conducting covert attacks on walking victims in outdoor settings.
-In a real-world scenario illustrated in Fig.
-13b, we recruited volunteers carrying various types of devices (specifically for barebones testing) for practical validation at different walking speeds.
-The volunteers with victimized devices maintained walking speeds of approximately 1 m/s, 1.5 m/s, corresponding to normal walking and brisk walking states, respectively.
-The experimental outcomes depicted in Fig.
-20 demonstrate that when a victim moved at a speed of 1 m/s (average human walking pace) and was 5 m away from the MetaAttack, the success rate remained at 51.4%.
-It was observed that while MetaAttack maintains a consistent attack distance, the success rate of the attack significantly drops as the victim moves at a faster pace.
-Moreover, the efficacy of the MetaAttack varies across different victim devices.
-Related research [55, 56, 57] indicates that a distance of 3-4 m between individuals signifies a public distance suitable for interactions with unfamiliar individuals.
-Consequently, we believe that under regular walking circumstances, a trailing attack using MetaAttack beyond the three-meter threshold is unlikely to alert the victim, facilitating ease of success.
-To confirm the feasibility of MetaAttack, we released attack video against walking victims [32].
-Figure 20: The impact of moving target’s speeds (target’s speeds are 1 m/s and 1.5 m/s)
-VI-C3 S3 - Eavesdropping attacks in indoor environments
-Eavesdropping is a prevalent method of information theft in commercial activities, involving covertly monitoring private conversations or communications to obtain sensitive or confidential information for unauthorized purposes.
-In S3 (as shown in Figure 13c), attackers can accomplish eavesdropping by sending inaudible commands to the victim’s phone to initiate a call to a number pre-set by the attacker.
-By leveraging the victim devices for eavesdropping, there is no necessity to physically retrieve the eavesdropping device from the premises, thereby mitigating the risk of exposure.
-Nonetheless, the primary challenge in this scenario lies in necessitating enhancements in character recognition accuracy.
-Therefore, we conducted tests to measure the word accuracy of MetaAttack in transmitting the command “call 1234567890” to various victims’ devices across different distances.
-As depicted in Fig.
-24, the experiments indicated that MetaAttack achieved 100% word accuracy when targeting an Apple iPhone 14 Pro at a set attack distance of 8.7 m.
-For other devices, MetaAttack achieved 100% word accuracy within an average attack range of 8 m, demonstrating its capability to conduct eavesdropping attacks in most indoor settings.
+14a, the experiment in Scenario S1 was conducted between two high-rise buildings, taking into account high-altitude wind interference.
+The attacker was located in a public corridor with no power supply and frequent pedestrian traffic, making it impractical to use large equipment.
+The experimental results (Fig.
+20) show that MetaAttack maintained a stable attack range of 8.6 meters at wind speeds of 2–3 m/s, while wind speeds above 3.5 m/s slightly affected the attack distance.
+However, attackers can rely on weather forecasts to avoid unfavorable wind conditions and ensure reliable long-range attacks.
+VI-C2 S2 - Attack against outdoor targets
+Due to its portability, MetaAttack is well-suited for conducting covert attacks on moving targets in outdoor environments.
+To validate this, we recruited volunteers carrying different types of devices in the real-world scenario shown in Fig.
+14b.
+The volunteers walked at speeds of 1 - 1.5 m/s, corresponding to normal walking and brisk walking, respectively, with the devices either held in hand or placed in pockets.
+The experimental results in Fig.
+22 show that when the target moves at a speed of 1 m/s and holds the device, MetaAttack achieves its optimal average RSA of 6.2 meters.
+As the target’s speed increases or the device is placed in a pocket, the attack success rate decreases.
+This is due to slight shaking of the device during movement, making it difficult for the system to align, and physical obstructions like a pocket weakening the signal.
+However, even in the harshest conditions (1.5 m/s speed with the device in the pocket), MetaAttack still achieves an average RSA of 4.1 meters.
+As mentioned in Section II-B, a distance beyond 3 meters is considered a safe range unlikely to alert the target.
+Therefore, we believe that MetaAttack can perform attacks without alerting the target.
+To validate the feasibility of MetaAttack, we released an attack video targeting walking individuals [37].
+Figure 21: Impact of outdoor target.
+Figure 22: System’s Word Acc. (WA)
+VI-C3 S3 - Eavesdropping attacks on sensitive information
+Eavesdropping is a common method of information theft in commercial activities, where attackers send inaudible commands to manipulate the target phone to dial a preset number, as shown in Fig.
+14c.
+The main challenge in this scenario is improving character recognition accuracy.
+We evaluated the performance of MetaAttack in transmitting the command call 1234567890 at various distances.
+Results show that MetaAttack achieved 100% recognition accuracy at a distance of 8.85 meters on an Apple iPhone 14 Pro, and similar accuracy on other devices within an average range of 8 meters.
+This demonstrates that the metamaterial-based beamforming design of MetaAttack enables efficient and precise signal transmission.
 VI-C4 S4 - Effect of passing individuals on attack range
-Passing individuals can significantly impact the propagation of inaudible sound waves.
-To investigate the influence of obstacles on the range of MetaAttack attacks at various distances, we conducted experiments using a public study room setting as an illustrative scenario.
-The detailed experimental parameter settings are shown in Fig.
-13d.
-The experiment involved two passing individuals.
+Passing pedestrians may affect the propagation of inaudible sound waves.
+To assess their impact on the attack range of MetaAttack, we conducted experiments in a public study room (parameters shown in Fig.
+14d).
 As shown in Fig.
-23, taking the iPhone 14 Pro as an example, when someone crosses MetaAttack’s attack path back and forth at a speed of 1 m/s, the system’s performance is minimally affected, allowing the attack to reach average distances of up to 7.6 m.
-Experimental findings indicate that passing individuals absorb and scatter ultrasonic waves, gradually attenuating the transmitted waves.
-But it will not significantly affect the attack range of MetaAttack.
-Figure 21: Effect on attack range by passing people
-Figure 22: Attack Range vs. Number of Speakers
-Figure 23: The feedback accuracy of MetaAttack for different scenarios and devices
-VI-C5 S5 - Feedback on the results of an inaudible attack
-The feedback mechanism is crucial for inaudible attacks, as it enables the attacker to swiftly clean the attack site to evade detection by the victim.
+24, when a person walks back and forth along the attack path at a speed of 1 m/s (using the iPhone 14 Pro as an example), the attack distance still reaches 7.6 meters with almost no performance degradation.
+Although pedestrians may absorb and scatter part of the ultrasonic signal, the beamforming design of MetaAttack ensures that the sound waves remain highly directional and focused, allowing the signal to be transmitted effectively without being significantly dispersed, thus maintaining the effectiveness of the attack.
+Figure 23: Effect by passing people
+Figure 24: Performance compared to prior research (attack range vs. number of speakers)
+Figure 25: The feedback accuracy of MetaAttack for different scenarios
+VI-C5 S5 - Feedback mechanism performance
+The feedback mechanism is crucial for inaudible attacks, as it enables the attacker to swiftly clean the attack site to evade detection by the target.
 In this experiment, we evaluated the efficacy of MetaAttack’s feedback across scenarios S1 to S4.
-The experimental results, depicted in Fig.
-23, demonstrate varying Precision and Recall values across different devices.
-Notably, MetaAttack exhibited enhanced performance when victim devices were Apple.
-Conversely, MetaAttack exhibits poorer performance in unstable scenarios (Figures 13a and 13b), indicating that external factors have a greater impact on the feedback mechanism.
-In such instances, the attacker must exercise timely judgment based on personal experience to avoid detection by the victim, particularly in the event of a successful and continued attack.
-Figure 24: The word accuracy of MetaAttack for different devices
+As shown in Fig.
+25, although the Precision and Recall vary slightly across different devices, they consistently remain above 97%, demonstrating high feedback accuracy.
+Notably, MetaAttack performs relatively worse in unstable scenarios (Fig.
+14a and Fig.
+14b), indicating that the feedback mechanism is more susceptible to external environmental interference.
+In such instances, the attacker must exercise timely judgment based on personal experience to avoid detection by the target, particularly in the event of a successful and continued attack.
 VI-D Performance Compared to Prior Research
@@
 As shown in Fig.
-23, the MetaAttack uses only 12 speakers for its attacks and relies on only two 70 mm $\times$ 130 mm lithium batteries for power.
+24, the MetaAttack uses only 12 speakers for its attacks and relies on only two 70 mm $\times$ 130 mm lithium batteries for power.
 LipRead [19] used 61 speakers for its operations and relied on power from a DC power supply.
@@
 Hardware devices for inaudible attacks.
-Investigating the influence of distinct attack devices on the functionality of MetaAttack system represents an interesting direction for future research.
-Through experimentation with a range of speaker and amplifier models, it is expected that MetaAttack system may potentially enhance its attack range efficiency while optimizing resource consumption.
-This aspect is worthy of further exploration and analysis in future research endeavors.
+Exploring the impact of different attack devices on the functionality of the MetaAttack system is an interesting direction for future research.
+By experimenting with various speaker and amplifier models, the MetaAttack system is expected to improve attack range efficiency while optimizing resource consumption.
 Attack through walls, windows, or curtains.
 The attenuation of ultrasonic signals when passing through solid objects poses a significant challenge for transmitting inaudible commands at 40 kHz.
-This makes it difficult for MetaAttack to effectively attack victims in enclosed environments.
+This makes it difficult for MetaAttack to effectively attack targets in enclosed environments.
 Future work will focus on addressing this issue.
+Audible leakage in the air propagation process.
+The audible leakage during airborne transmission caused by nonlinear acoustics remains an issue to be addressed.
+Although MetaAttack has compressed the audible leakage into a narrow range of about 12°, where it can only be detected if the target’s ear is within this range, the problem is not yet fully resolved.
+We plan to thoroughly address this limitation in future work.
+Optimization of Metamaterials.
+Currently, MetaAttack has achieved the longest possible attack range under its existing structure.
+We believe that further extending the attack distance hinges on enhancing the filtering performance of the metamaterials and improving their beamforming capability at lower carrier frequencies (e.g., around 25 kHz).
+Compared to high-frequency ultrasound, low-frequency ultrasound maintains inaudibility while offering better propagation characteristics, making it more promising for long-range transmission.
+In addition, due to the constraints of the fixed LAM filtering structure, the current optimization of speaker parameters (such as quantity, arrangement, and carrier frequency) may only yield locally optimal results.
+In future work, we plan to explore more advanced metamaterial designs and jointly optimize them with speaker parameters to enable more stable and longer-range attacks.
 Possible conflicts in the feedback mechanism.
-Considering that a few attack commands (such as ”turn on airplane mode”) may conflict with feedback mechanism, we can first confirm the position of the victim device through feedback mechanism before executing the attack.
+Considering that a few attack commands (such as ”turn on airplane mode”) may conflict with feedback mechanism, we can first confirm the position of the target device through feedback mechanism before executing the attack.
 VIII COUNTERMEASURES
 Previous defense strategies.
-The previous defense strategies against inaudible attacks can be mainly divided into two categories: forensics-based defense strategies and cancellation-based defense strategies.
-Forensics-based defense strategies utilize the nonlinear characteristics of inaudible commands to identify them by detecting specific differences between these commands and legitimate commands [8, 14].
-In contrast, cancellation-based defense strategies by using additional ultrasonic speaker arrays to emit sound waves that cancel out the inaudible commands [58].
+Defense strategies against inaudible attacks fall into two main categories: forensics-based and cancellation-based.
+Forensics-based defenses detect attacks by analyzing signal characteristics through methods like frequency analysis or deep learning.
+For example, EarArray [62] analyzes timing differences across microphones; Li et al.
+[65] extract spatial features and apply deep learning to distinguish human speech from machine-generated signals; NormDetect [63] identifies missing features in attack audio; and LipRead [19] analyzes power, autocorrelation, and low-frequency amplitude deviations.
+Cancellation-based defenses use ultrasonic speaker arrays to emit sound waves that neutralize attack signals [66].
 Challenges and future directions in defending against MetaAttack attacks.
-Forensics-based defense strategies can only issue warnings upon detecting an attack, while attacks from MetaAttack can often succeed within a few seconds, making it difficult for victims to timely stop ongoing attacks [58].
-Although cancellation-based defense strategies can effectively prevent ongoing attacks, their effective range is limited to just a few meters.
-Thus, when defending MetaAttack that can conduct outdoor mobile attacks, victims must carry cumbersome equipment, such as speaker arrays and laptops, which increases inconvenience.
-We believe that future research should focus on developing more portable protective devices that can be carried easily or directly installed on smart devices to address the threats posed by MetaAttack.
+Forensics-based defenses are sensitive to microphone variations, leading to unstable results [66].
+We replicated LipRead [19] in a quiet (40 dB) lab and tested 30 rounds on various devices.
+iPhone 13 Pro Max, iPhone 13, iPhone 12, Galaxy S7, and nova 7 SE 5G showed over 90% success, while iPhone 14 Pro, Redmi K50 Ultra, Pixel 3, and Pixel 2 were below 63.3%, confirming poor cross-device consistency.
+These defenses also disable voice assistants only after detecting attacks, disrupting normal use.
+Cancellation-based methods are more effective but rely on bulky equipment and have limited range, making them impractical for mobile attacks like MetaAttack.
+Future defenses should focus on lightweight, externally integrated mechanisms that block attack signals before they reach the microphone.
 IX RELATED WORK
-IX-A Attacks on Voice Control Systems
-Attacks on voice control systems have been proven to cause serious security issues.
-The previous attacks are mainly into two types: attacks based on adversarial examples and inaudible attacks.
-The attacks based on adversarial examples are presented by deceiving the victims [59, 60, 61].
-However, these attacks based on adversarial examples rely on audible sounds, which can cause suspicion among alert victims.
-Inaudible attacks can directly disrupt or control voice control systems through ultrasound, which makes them more covert [7, 8, 9, 14].
-For example, DolphinAttack [8] modulates attack commands onto ultrasound waves for attack, achieving control over smart devices.
-LipRead [9] and DolphinAttack (Long-range System) [15] expands the range of inaudible attacks on this basis, but its large devices are easily detected.
-Our MetaAttack first combines attacks with acoustic metamaterials to achieve long-range attacks in a compact size, making it more covert compared to previous studies.
-IX-B Acoustic Metamaterials
-Acoustic metamaterial designed for ventilation barriers is an emerging research field.
-This acoustic metamaterial can achieve the filtering function while maintaining ventilation through compact size.
+Attacks on Voice Control Systems.
+Attacks on voice control systems have been proven to pose serious security threats.
+Existing attack methods mainly fall into the following categories: adversarial attacks, magnetic injection attacks, laser attacks, and inaudible attacks.
+Adversarial attacks deceive devices by embedding carefully crafted instructions in normal audio to induce them to perform actions preset by the attacker [67, 68, 69, 70, 71, 72, 73, 74, 75].
+For example, CommanderSong [67] hides commands within songs and performs the attack by playing the song.
+Metamorph [68] enhances robustness using pre-coding, extending the attack range to 6 meters.
+In addition, MagBackdoor [17] injects command signals into voice control systems through the device’s charging port, while Sugawara et al.
+[76] use long-range lasers to simulate sound vibrations to attack voice control systems.
+Inaudible attacks can silently interfere with or control voice control systems, making them more covert [11, 12, 13, 19, 42, 20, 18, 46].
+For example, DolphinAttack [12] achieves control over smart devices by modulating attack commands onto ultrasound waves.
+LipRead [13] and DolphinAttack (Long-range System) [20] further extend the attack range by incorporating spectrum splitting techniques and large speaker arrays.
+VRIFLE [18] injects ultrasonic signals while the target is speaking, interfering with the target’s control of the device, thereby increasing the destructiveness of the attack.
+MetaAttack first combines attacks with acoustic metamaterials to achieve long-range attacks in a compact size, offering both concealment and threat.
+Acoustic Metamaterials.
+Acoustic metamaterials designed for ventilation barriers are an emerging research area, enabling both filtering and airflow in a compact form.
 For example, Ghaffarivardavagh et al.
-[42] proposed a deep subwavelength acoustic metamaterial using Fano-like resonance.
-Subsequently, Zhu et al.
-[25] increased the filtering bandwidth of the acoustic metamaterial through nonlocal effects (600 Hz - 1900 Hz).
-Based on the above research, MetaAttack further expands the filtering range of this type of metamaterials (100 Hz - 4000 Hz) and first expands its application field to the inaudible attacks.
+[51] proposed a deep subwavelength acoustic metamaterial based on Fano-like resonance, and Zhu et al.
+[30] expanded its filtering bandwidth using nonlocal effects.
+Building on this, MetaAttack further extends the filtering range to 100 – 4000‘Hz and pioneers its application in ultrasonic attacks.
 X CONCLUSION
@@
 Our approach can achieve covert and longer-range inaudible attacks using a smallest device compared to prior works.
+XI Ethical Considerations
+Ethical Challenges in Reproducibility and Disclosure.
+Although publishing the 3D model designs of MetaAttack would aid in the reproducibility of the experiment, it could also provide malicious actors with the information necessary to execute attacks.
+Therefore, we have decided to share the related materials only with verified researchers who contact us, balancing transparency with security.
+Ethical Considerations in Experiments and Testing.
+The research equipment was purchased at our own expense, and all experiments were approved by the university’s Institutional Review Board (IRB), ensuring compliance with data security and ethical standards.
+All participants voluntarily agreed to participate in the experiments based on informed consent, ensuring adherence to ethical requirements.